### v2

# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

Abstract
The disordered quantum systems host three types of quantum states, the extended, localized, and critical, which bring up various distinct fundamental phases, including the pure phases and coexisting ones with mobility edges. The quantum phases involving critical states are of particular importance, but are less understood compared with the other ones, and the different phases have been separately studied in different quasiperiodic models. Here we propose a unified framework based on a spinful quasiperiodic system which unifies the realizations of all the fundamental Anderson phases, with the exact and universal results being obtained for these distinct phases. Through the duality transformation and renormalization group method, we show that the pure phases are obtained when the (emergent) chiral symmetry preserves in the proposed spin-1/2 quasiperiodic model, which provides a criterion for the emergence of the pure phases or the coexisting ones with mobility edges. Further, we uncover a new universal mechanism for the critical states that the emergence of such states is protected by the generalized incommensurate matrix element zeros in the spinful quasiperiodic model, as a novel generalization of the quasiperiodic hopping zeros in the spinless systems. We also show with the Avila’s global theory the criteria of exact solvability for the present unified quasiperiodic system, with which we identify several new quasiperiodic models derived from the spinful system hosting exactly solvable Anderson phases. In particular, we reach a single model that hosts all the seven fundamental phases of Anderson localization. Finally, an experimental scheme is proposed to realize these models using quasiperiodic optical Raman lattices. This study establishes a unified theoretical framework which enables an in-depth exploration of the broad classes of all fundamental localization phases in quasiperiodic systems, and offers key insights for constructing their exactly solvable models with experimental feasibility.
I Introduction
Anderson localization, wherein quantum states localize exponentially in real space due to disorder-induced scattering, is a fundamental and universal phenomenon in disordered systems (Anderson, 1958; Lee and Ramakrishnan, 1985; Kramer and MacKinnon, 1993; Evers and Mirlin, 2008). It manifests distinctly across two categories of aperiodic systems: disordered systems and quasiperiodic lattices. In disordered systems, scaling theory predicts that noninteracting states are localized in one and two dimensions even under weak disorder, whereas in three dimensions, an Anderson transition from extended to localized states occurs only with sufficiently strong disorder (Abrahams et al., 1979). In contrast, quasiperiodic (QP) lattices exhibit richer physics including Anderson transitions in all spatial dimensions, governed by quasiperiodic parameters (Aubry and André, 1980; Roati et al., 2008). Importantly, both extended and localized states can coexist within a single quantum phase, separated by characteristic energies known as mobility edges (MEs) (Soukoulis and Economou, 1982; Das Sarma et al., 1988; Boers et al., 2007; Biddle and Das Sarma, 2010; Ganeshan et al., 2015; Li et al., 2017; Gopalakrishnan, 2017; An et al., 2018; Lüschen et al., 2018; Kohlert et al., 2019; Deng et al., 2019; Wang et al., 2020a; Li and Das Sarma, 2020; Duthie et al., 2021; An et al., 2021; Roy et al., 2021; Roy and Sharma, 2021; Wang et al., 2021a, 2022a; Longhi, 2023; Wang et al., 2023; Gao et al., 2023; Longhi, 2024a, b; Wei et al., 2024; Gao et al., 2024; Liu et al., 2024; Tabanelli et al., 2024; Shimasaki et al., 2024a). The ability to derive exact analytical expressions for characteristic length sclaes and MEs in various QP models Wang et al. (2020a, 2021a); Zhou et al. (2023); Gonçalves et al. (2023a); Liu et al. (2022) plays a crucial role in advancing the fundamental understanding of localization transitions, providing theoretical insights that extend beyond numerical simulation.
Between the extended and localized states lie the critical states, which are delocalized in both position and momentum spaces. Critical states have recently attracted extensive research interests (Hatsugai and Kohmoto, 1990; Han et al., 1994; Tanese et al., 2014; Zhou et al., 2023; Wang et al., 2016; Liu et al., 2015; Zeng et al., 2016; Yao et al., 2019; Wang et al., 2020b; Goblot et al., 2020; Xiao et al., 2021; Liu et al., 2022; Li et al., 2023; Wang et al., 2022b; Gonçalves et al., 2023a; Lin et al., 2023; Lee et al., 2023; Dai et al., 2023; Shimasaki et al., 2024b; Li et al., 2024; Wang and Zhou, 2025; Huang et al., 2025; Chen et al., 2024; Guo, 2024; Dotti et al., 2024; Yang et al., 2024; Duncan, 2024; Yao et al., 2024; Bai and Weld, 2025) due to their unique properties, such as local scale invariance, multifractality, and critical quantum dynamics. Recent advances, particularly through Avila’s global theory (Avila, 2015; Zhou et al., 2023), refine a rigorous characterization Simon and Spencer (1989); Jitomirskaya and Marx (2012) of the critical states, showing that critical states generally arise when the hopping couplings in a QP system are non-uniform and possess incommensurately distributed zeros (IDZs) in the thermodynamic limit Liu (2024). The exact theory enables unambiguous determination of critical phases and the new MEs between critical and other states Zhou et al. (2023). Furthermore, this rigorous characterization of critical states provides a well-defined foundation for incorporating many-body interactions, thereby extending the paradigm of criticality beyond the single-particle picture. Inclusion of interactions further enriches the physics of the critical phase, with multifractality of the wave function influencing both ground state properties associated with symmetry breaking (Liu, 2024; Feigel’man et al., 2007; Burmistrov et al., 2012; Zhao et al., 2019; Sacépé et al., 2020; Gonçalves et al., 2024) and the emergence of many-body critical (MBC) phases (Wang et al., 2021b, 2020b) at infinite temperature. The MBC phase, interpolating the thermal phase and many-body localization (MBL) (Pal and Huse, 2010; Nandkishore and Huse, 2015; Schreiber et al., 2015; Bordia et al., 2016, 2017a; Lüschen et al., 2017; Bordia et al., 2017b), signifies a non-ergodic but delocalized regime that violates the eigenstate thermalization hypothesis (ETH) (D’Alessio et al., 2016; Deutsch, 1991; Rigol et al., 2008; Srednicki, 1994).
The recently developed quasiperiodic mosaic lattice models (Wang et al., 2020a; Zhou et al., 2023; Gao et al., 2024; Huang et al., 2025) and their generalizations Longhi (2023); Dai et al. (2023); Lin and Gong (2024); Guo (2024); Wei et al. (2024); Longhi (2024a) advance the rigorous characterization of all eigenstates in the spectra through the analytical study of their localization properties (Lyapunov exponent or LE) using Avila’s global theory (Avila, 2015). With the mosaic models, the MEs can also be analytically obtained. On the other hand, while the various exactly solvable models are proposed, these different models are treated on a case-by-case basis. There is no unified theory showing that under what conditions a generic exactly solvable model undergoes phase transitions between the pure localized, extended, and critical phases without MEs, and otherwise it possesses MEs that separate these states, i.e., the transitions are energy-dependent. Besides, Avila’s global theory enables an analytic study of the spinless models with nearest-neighbour hopping, while the exact results for more complicated systems are scarce, hindering the in-depth understanding of the localization physics in more generic systems. For instance, the universal mechanism for critical states in the complex systems with internal degrees of freedom, such as spins, remains unknown. Addressing these issues is not only important in fundamental theory, but also of high significance in experiment, as the remarkable experimental progresses in engineering the spin and orbital degrees of freedom Zhang and Liu (2018); Liu et al. (2014); Wu et al. (2016); Wang et al. (2018); Sun et al. (2018); Song et al. (2019); Lu et al. (2020); Wang et al. (2021c); Zhang et al. (2024) have enabled the realization of complex QP systems beyond toy models. These crucial fundamental issues motivate us to establish a universal theory that can generically characterize spinful QP lattice systems without relying on specific microscopic details.
Figure 1: Illustration of generic framework. (a) The different processes of the spinful quasiperiodic systems. The hopping coupling matrix $\Pi_{j}^{s,s^{\prime}}$ denote the momentum transfer coupling with the diagonal (off-diagonal) terms for the same (different) internal degrees of freedom. The on-site matrix $M_{j}^{s,s^{\prime}}$ represents the on-site spin flipped and on-site modulation of the system. (b) The conditions when the Avila’s global theory can give rise to the analytical characterization of the (reduced) 1D quasiperiodic chain. The nonzero acceleration of the complexified Lyapunov exponent $\gamma_{\epsilon}$ is quantized to the unique integer: $d\gamma_{\epsilon}/d\epsilon=2\pi\mathbb{Z}$. Then the original Lyapunov exponent can be obtained by the expression obtained from the $\epsilon\rightarrow\infty$.
I.1 Summary of results
In this work, we propose to investigate the generic spin-$1/2$ quasiperiodic (QP) system, with which we establish a unified theoretical framework for all fundamental localization phases based on the exact and universal results uncovered in this system. The proposed generic QP framework not only unifies existing 1D spinful QP models with exact analytical solutions, but also encompasses spinless QP models through an additional Majorana representation. More importantly, this system establishes a versatile platform for constructing new exactly solvable QP models exhibiting rich phase structures, including pure phases and coexisting phases with MEs.
The universal results are obtained in this unified framework and can be summarized as follows. First, we show with duality transformation and renormalization group method that the pure extended, localized, or critical phases are obtained when the system preserves a chiral symmetry. Accordingly, the emergence of coexisting phases with MEs necessitates to break the chiral symmetry. Second, we unveil a new universal mechanism for critical states in the spinful quasiperiodic system, which arise from generalized incommensurate zeros (GIZs) in matrix elements. This mechanism is a novel generalization of the incommensurate hopping zeros in the spinless QP models, enabling a much broader pathway of realizing and characterizing the critical states. Third, we establish the exactly solvable condition, under which the 1D spinful QP system can be transformed into effectively spinless QP models of dressed particles, and then can be exactly solved. This result provides a powerful guiding principle to construct exactly solvable models that host a variety of novel localization phases with or without MEs, particularly the models encompassing all seven fundamental phases in localization physics.
I.2 Organization of the work
The remainder of the paper is organized as follows. In Sec. II, we introduce the generic framework for 1D spinful QP chains, including the basic Hamiltonian and the analytical approaches employed in this study, such as dual transformations, renormalization group methods, and Avila’s global theory. In Sec. III, we present the universal results for the spinful QP chains, as organized into three theorems shown here. These universal results serve as guiding principles for constructing multiple exactly solvable models hosting distinct classes of localization physics and anomalous MEs, including a novel model that realizes the seven fundamental phases of Anderson localization in Sec. IV. In Sec. V, we propose an experimental scheme for the realization of these new models using a generic quasiperiodic optical Raman lattice. Finally, we conclude in Sec. VI, with additional details and proofs provided in Appendix.
II Generic Framework
II.1 The spinful quasiperiodic model
We begin with the generic one-dimensional (1D) spin-$1/2$ quasiperiodic system, whose Hamiltonian is given by
$H=\sum_{j,s,s^{\prime}}\big{(}c_{j+1,s}^{\dagger}\Pi_{j}^{s.s^{\prime}}c_{j,s^{\prime}}+\mathrm{h.c.}\big{)}+\sum_{j,s,s^{\prime}}c_{j,s}^{\dagger}M_{j}^{s,s^{\prime}}c_{j,s^{\prime}},$ (1)
where $c_{j,s}^{\dagger}(c_{j,s})$ creates (annihilates) a particle on site $j$ with (pseudo)spin $s=\uparrow,\downarrow$ that can denote real spin, sublattices or particle-hole. This model includes two key ingredients, as illustrated in Fig. 1(a). First, the hopping coupling matrix $\Pi_{j}=\sum_{l}p_{j}^{l}\sigma_{l}$, as written in terms of Pauli matrices $\sigma_{l}$ with $l=\{0,x,y,z\}$, represents the spin-independent (for $l=0$) and the spin-dependent (for $l=x,y,z$) hopping coupling terms. The latter includes both the spin-conserved ($l=z$) and spin-flip ($l=x,y$) hopping terms, corresponding to the spin-orbit coupling in the periodic systems. Second, the onsite matrix $M_{j}=\sum_{l}m_{j}^{l}\sigma_{l}$ also involves the spin-independent chemical potential and the Zeeman coupling terms. The coefficients $p_{j}^{l}=t^{l}+\mu^{l}V_{j}^{\mathrm{od}}$ and $m_{j}^{l}=\lambda^{l}+V^{l}V_{j}^{\mathrm{d}}$, where $(t^{l},\lambda^{l},\mu^{l},V^{l})$ are constants, the off-diagonal quasiperiodic modulation $V_{j}^{\mathrm{od}}=\cos[2\pi\alpha(j+1/2)+k_{y}]$ and the diagonal quasiperiodic modulation $V_{j}^{\mathrm{d}}=\cos(2\pi\alpha j+k_{y})$. Here $\alpha=\lim_{n\rightarrow\infty}(F_{n-1}/F_{n})=(\sqrt{5}-1)/2$ is an irrational number approached by the Fibonacci sequence $F_{n}$, and $k_{y}$ is a constant phase factor. In this work, we take that the quasiperiodic term has a single frequency. To facilitate our discussions, we define the combinations of the Pauli matrices $\sigma_{l}$ as $\sigma_{\pm}=(\sigma_{x}\pm+i\sigma_{y})/2$ and $\Lambda_{\pm}=(\sigma_{z}\pm\sigma_{0})/2$.
This spinful QP system unifies various QP models, including the existing 1D spinful models with analytical solutions, and spinless models by taking the Majorana representation (see the details in Appendix C and Appendix D). More importantly, with this unified framework we shall develop a series of new and exactly solvable models which host various novel quantum phases with or without mobility edges. Below, we outline the three main theoretical approaches employed in this study.
II.2 Dual transformation
The dual transformation is a non-local operation similar to the Fourier transform, and maps the system between real space and dual momentum space. It is defined as $c_{j,s}=(1/\sqrt{L})\sum_{n}e^{-i2\pi n\alpha j}c_{n,s}$. The extended, localized, and critical states behave differently under this transformation. The critical states exhibit behavior distinct from the other two. For an extended state, the wave function is delocalized in real space, but localized in dual momentum space, whereas the localized state is delocalized in dual momentum space but localized in real space. These opposing features are exchanged under the dual transformation. However, critical states remain delocalized in both real and dual spaces and are invariant under the transformation. For example, when the system has a QP hopping term $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$, which dominates over other hopping couplings and is invariant under the dual transformation due to the presence of incommensurately distributed zeros (IDZs) in the QP hopping modulation $V_{j}^{\rm od}$, the critical states may emerge and exhibit self-dual for having incommensurate nodes corresponding to the IDZs Zhou et al. (2023). These incommensurate nodes in the delocalized wave functions remain unchanged under the dual transformation, driving the delocalized states into critical states. An intuitive picture suggests that IDZs partition the 1D system into multiple segments, where the wave function reorganizes, exhibiting a self-similar pattern characteristic of critical states.
The distinct behaviors of the three types of states can be mapped onto a 2D system, providing an intuitive physical interpretation, where $x$- and $y$-axes correspond to real and dual space, and dual transfomration coppresonds to an exchange between two axes. In this mapping, the extended state is delocalized along $x$-direction and localized along $y$-direction, and the opposite is true for localized states. The critical state is delocalized in both directions. Thus, the swap of the two axes transforms the localized and extended states, while leaving the critical states unchanged. We use 1D spinless 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) as an example, with the nearest neighbor hopping $t_{j}=t+\mu V_{j}^{\rm od}$ and on-site potential $V_{j}=V_{0}V_{j}^{\rm d}$ (see details in Appendix D). By interpreting the phase shift $k_{y}$ as the quasi-momentum in the $y$-direction and performing a Fourier transform, the 1D spinless EAA model maps to a 2D Hofstadter model that describes a particle moving in a magnetic field (Kraus and Zilberberg, 2012; Borgnia and Slager, 2023). In this mapping, the uniform hopping terms correspond to hopping in the $x$-direction, while the on-site potential $V_{j}^{\mathrm{d}}a_{j}^{\dagger}a_{j}$ corresponds to the hopping in the $y$-direction with a magnetic flux $\exp(i2\pi\alpha x)a_{x,y}^{\dagger}a_{x,y+1}+{\rm h.c.}$. Similarly, the QP hopping terms $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$ represent diagonal hopping in the 2D plane under the magnetic flux. The relative hopping amplitude shapes the cyclotron orbits while maintaining the flux within each orbit. Dominant hopping in the $x$- and $y$- directions in 2D facilitates the emergence of extended, localized states. In contrast, dominant diagonal hopping, giving equal weight to both directions, results in critical states.
II.3 Renormalization group
The renormalization group (RG) approach (Suslov, 1982; Thouless and Niu, 1983; Ostlund and Pandit, 1984; Niu and Nori, 1986, 1990; Szabó and Schneider, 2018; Jagannathan, 2021; Gonçalves et al., 2022, 2023b) investigates the relevance of renormalized coefficients by iterating the rational approximations of the QP parameter. Through this process, the RG determines the asymptotic behavior of the target eigenstates. With the rational approximation, the approximated Hamiltonian exhibits a periodic structure, where the energy dispersion becomes a periodic function of quasi-momentum $\kappa_{x}$ and $\kappa_{y}$. To systematically investigate how the dispersion coefficients evolve as the system size increases, we compute the characteristic determinant $P^{(n)}=|{\cal H}^{(n)}-E|$ at a target energy $E$ for a given $L=F_{n}$, which is given by
$P^{(n)}(E;\kappa_{x},\kappa_{y})$ $=t_{R}^{(n)}\cos(\kappa_{x}+\kappa_{x}^{0})+V_{R}^{(n)}\cos(\kappa_{y}+\kappa_{y}^{0})$
$+\mu_{R}^{(n)}\cos(\kappa_{x}+\tilde{\kappa}_{x}^{0})\cos(\kappa_{y}+\tilde{\kappa}_{y}^{0})$
$+\epsilon_{R}^{(n)}(E,\varphi,\kappa)+T_{R}^{(n)}(E),$ (2)
where $\epsilon_{R}^{(n)}$ represents higher harmonic terms. The asymptotic nature of eigenstates at $n\rightarrow\infty$ is determined by comparing the relative magnitudes of renormalized coefficients. For extended states, the dominant hopping is along $x$ direction, yielding $\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|},\big{|}V_{R}^{(n)}/t_{R}^{(n)}\big{|}\rightarrow 0$. For localized states, the dominant hopping is along $y$ direction, leading to $\big{|}t_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|}\rightarrow 0$. For critical states, all the three coefficients reamin comparable, such that $\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|}\geq 1$. The corresponding transition points between extend, localized and critical can be obtained accordingly, and the mobility edges corresponds to the energy dependent transition points.
II.4 Avila’s global theory
When the Hamiltonian in Eq. (1) can be reduced to an effectively spinless quasiperiodic (QP) chain of dressed particles, the system can be analytically characterized using Avila’s global theory Avila (2015). This theory enables the analytical calculation of the Lyapunov exponent (LE), which is the inverse of the localization length. The method provides a rigorous characterization of extended, localized, and critical states. Specifically, the non-negative LE $\gamma(E)$, quantifies the localization properties of eigenstates with energy $E$. If $\gamma(E)>0$, the state is localized, with the localization length $\xi(E)=\gamma^{-1}(E)$. If $\gamma(E)=0$, the state is delocalized, with a divergent localization length, corresponding to either an extended or a critical state. Critical states arise when the transfer matrix $T_{j}$ of the delocalized states becomes singular Simon and Spencer (1989); Jitomirskaya and Marx (2012). This singularity originates from the presence of incommensurately distributed zeros (IDZs) in the hopping coefficients Zhou et al. (2023), or from an unbounded QP potential, which effectively divides the system into infinite segments, acting as IDZs.
Furthermore, Avila’s global theory can also be employed to derive the analytical expression for the correlation length of extended states. This is achieved by performing the dual transformation of the Hamiltonian into dual space, followed by calculating the localization length of localized states in dual space, which is equivalent to the correlation length of the corresponding extended states in real space.
In the following, we outline the procedure for obtaining the LE using Avila’s global theory. For the generic 1D spinful QP Hamiltonian in Eq. (1), we consider the 4-by-4 transfer matrix $T_{j}$ given by $(\begin{array}[]{cccc}u_{j+1\uparrow},&u_{j+1\downarrow},&u_{j\uparrow},&u_{j\downarrow}\end{array})^{\intercal}=T_{j}(\begin{array}[]{cccc}u_{j\uparrow},&u_{j\downarrow},&u_{j-1\uparrow},&u_{j-1\downarrow}\end{array})^{\intercal}$, where $u_{j\sigma}$ is the wave function of eigenstate at site $j$ with spin $\sigma$. However, the 4-by-4 transfer matrix does not guarantee exact solvability. Exact solvability arises when the transfer matrix is reduced to a block-diagonal 2-by-2 form. When the system incorporates the local constraint introduced in Sec. III.3, the Hamiltonian simplifies to a 1D spinless QP model of dressed particles, featuring a single-frequency potential and nearest-neighbor hopping. Consequently, the effective transfer matrix becomes a 2-by-2 matrix. In this regime, the Lyapunov exponent $\gamma_{\epsilon}(E)$ is computed as
[…]
Here, $E$ is the energy of the corresponding eigenstate, $||A||$ is the norm of the matrix A, i.e. the square root of the largest eigenvalue of $A^{\dagger}A$, ${\cal T}_{n,1}=\prod_{j=1}^{n}T_{j}$. The $\epsilon$ is the imaginary part of complexified quasiperiodic phase shift $\theta+i\epsilon$. The LE is generally challenging to compute analytically, even for the simple 2-by-2 transfer matrix $T_{j}$. Nevertheless, a key result from Avila’s global theory allows for the computation of the LE for all eigenstates: The LE $\gamma_{\epsilon}(E)$ is convex, continuous and piecewise linear with quantized right-derivative, given by
$\lim_{\epsilon\rightarrow 0^{+}}\frac{1}{2\pi\epsilon}[\gamma_{\epsilon}(E)-\gamma_{0}(E)]=\mathbb{Z}.$ (4)
The LE can be analytically obtained when the nonzero derivative is quantized to be exactly one integer for all range of $\epsilon$: $d\gamma_{\epsilon}/d\epsilon=2\pi\mathbb{Z}$. This condition implies the absence of turning points before which corresponds to there is no turning point before $\gamma_{\epsilon}$ intersects the $x-$ or $y-$axis, as illustrated in Fig. 1(b). In this case, the analytic expression for the LE is first obtained by calculating the complexified LE $\gamma_{\epsilon}$ at the limit $\epsilon\to\infty$, which is computationally straightforward. Using the properties that $\gamma_{\epsilon}(E)$ is convex, continuous, and piecewise linear with a quantized right-derivative, the analytic expression of the LE determined at $\epsilon\to\infty$ can be extended to all values of $\epsilon$. Finally, the exact LE is obtained by evaluating $\gamma_{\epsilon=0}$, which represents the original Lyapunov exponent that characterizes the properties of all eigenstates.
Figure 2: Three universal results for the spinful quasiperiodic chains. (a) Criteria for the system to exhibit pure phases without mobility edges. Here, $\lambda$ is the system parameter and $E$ denotes the energy. A and B refer to two distinct phases, which may correspond to extended, localized, or critical states. The system transitions to pure phases in the absence of mobility edges, which implies that the energy-dependent transition points $\lambda_{c}(E)$ become energy-independent, denoted as $\lambda_{c}$. (b) Mechanism for the emergence of critical states in the 1D spinful quasiperiodic (QP) system, generated by dual-invariant generalized incommensurate zeros (GIZs) in matrix elements, marked as the blue and red circles in real and dual spaces. (c) The local constraint that transforms the generic 1D spinful QP chain into a 1D spinless QP chain of dressed particles (represented by orange spheres), with energy-dependent or independent nearest-neighbor hopping coefficients $t_{j}^{\mathrm{eff}}(E)$ and on-site potentials $V_{j}^{\mathrm{eff}}(E)$. This transformation can be analytically tracked using Avila’s global theory.
III Universal results for the quasiperiodic spinful chains
In this section, we show three universal results for the spinful quasiperiodic systems, which facilitates to establish the unified framework of all fundamental localization phases, as summarized in Fig. 2. First, we prove a generic condition under which mobility edges (MEs) disappear in the system, namely, the pure phases are obtained under this condition. Second, we uncover a new universal mechanism that critical states are protected by the generalized incommensurate zeros in matrix elements. Third, we outline the conditions, under which Avila’s global theory can be applied to obtain analytic solutions to the localization properties, correlation length, and the associated phase transition points and mobility edges.
III.1 Criteria for pure phases without MEs
For the system to exhibit a pure phase without MEs, it must satisfy the following criteria: The on-site matrix $M_{j}$ are purely quasiperiodic, and the hopping coupling matrix $\Pi_{j}$ are either uniform or purely quasiperiodic, with the system satisfying the chiral symmetry
$\sigma_{y}\Pi_{j}^{\prime}\sigma_{y}^{\dagger}=-\Pi_{j}^{\prime},\qquad\sigma_{y}M_{j}^{\prime}\sigma_{y}^{\dagger}=-M_{j}^{\prime}.$ (5)
Here $\Pi_{j}^{\prime}$ and $M_{j}^{\prime}$ are the transformed matrices obtained in the basis that diagonalizes the on-site coupling matrix, given explicitly as
$\Pi_{j}^{\prime}=p_{j}^{\perp}\sigma_{x}+p_{j}^{z}\sigma_{z},\qquad M_{j}^{\prime}=m_{j}^{z}\sigma_{z},$ (6)
We demonstrate this criterion by proving that all leading coefficients in Eq.(2) are energy-independent, ensuring eigenstate transitions occur independently of energy, and consequently, the system exhibits pure phases without mobility edges (MEs) [Fig.2(a)]. This follows from a careful power counting analysis of the energy dependence in characteristic polynomial $P(E;\kappa_{x},\kappa_{y})$ in Eq. (2). We consider an odd system size $L$ without loss of generality. Due to chiral symmetry, eigenenergies appear in pairs $(-E,E)$, and therefore $P(E)$ is an even function of energy $E$. In the power counting analysis, the energy $E$ explicitly contributes a power of $E$, while $p_{j}^{\perp}$ and $p_{j}^{z}$ contribute dispersions of $e^{ik_{x}}$, and $m_{j}$ contributes dispersions of $e^{ik_{y}}$. Here $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. We utilize the periodic structure to simplify the power counting. Thus, the characteristic polynomial can be expressed as the determinant of a $2L\times 2L$ matrix $|A(E)|$ (details in Appendix A), with elements
$A_{mj}=e^{i2\pi\alpha jm}\Big{[}m_{j}\sigma_{0}+a_{k_{x},k_{y}}^{11}\Lambda_{+}+a_{k_{x},k_{y}}^{22}\Lambda_{-}-E\sigma_{x}\Big{]},$ (7)
where
$a_{k_{x},k_{y}}^{11(22)}=t_{j-1}^{+(-)}e^{-i(2\pi\alpha m+k_{x})}+c.c.,$ (8)
with $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$. This structure ensures the energy terms and momentum dispersions decouple naturally, eliminating cross terms such as $Ee^{ik_{x}}$. Below, we separately analyze the dominant dispersions for extended, localized, and critical states.
For the extended and localized states, characterized by dispersions of the form $\cos n_{x}Lk_{x}$ and $\cos n_{y}Lk_{y}$, where $n_{x},n_{y}\geq 1\in\mathbb{Z}$, we show that the dominated coefficients are energy-independent. Specifically, the leading dispersion $\cos Lk_{x}$ and $\cos Lk_{y}$ vanish due to chiral symmetry, as these terms arise from the product of $L$ factors of $e^{ik_{x}}$ and $L$ factors of energy $E$, yielding an odd power of energy $E^{L}\cos Lk_{x}(k_{y})$, which is forbidden. Thus, the dominant terms become $\cos 2Lk_{x}$ and $\cos 2Lk_{y}$, whose coefficients are energy-independent, resulting solely from diagonal elements of the determinant, leaving no opportunity for involving $E$-dependent terms. Therefore, transitions associated with extended and localized states do not depend on energy, implying the absence of MEs.
For the critical states, characterized by dispersions of the form $\cos n_{x}Lk_{x}\cos n_{y}Lk_{y}$ with $n_{x},n_{y}\geq 1\in\mathbf{\mathbb{Z}}$, we similarly demonstrate energy independence. The chiral symmetry leads to two distinct scenarios: uniform or purely quasiperiodic hopping matrices $\Pi_{j}$, with $M_{j}$ always purely quasiperiodic (details in Appendix A).
In the first scenario, with unifrom $\Pi_{j}$, the leading dispersion $\cos Lk_{x}\cos Lk_{y}$ emerges from products of exactly $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$), making it energy-independent.
In the second scenario, where both $\Pi_{j}$ and $M_{j}$ are purely quasiperiodic, the leading dispersions include terms such as $\cos 2Lk_{x}\cos Lk_{y}$, $\cos Lk_{x}\cos 2Lk_{y}$, and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_dispersion , all of which are energy-independent. The dispersion $\cos Lk_{x}\cos 2Lk_{y}$ arises from product of $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$), containing exactly $2L$ factors, and thus remains free of energy-dependent contributions. Similarly, the other two dispersions originate from product of $2L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$), which generically lead to both $\cos 2Lk_{x}\cos Lk_{y}$ and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_rg , both of which are energy-independent. Consequently, critical states also exhibit energy-independent transitions, ensuring no associated MEs appear.
The proof of the criterion highlights the crucial role of chiral symmetry in excluding MEs for the present spin-$1/2$ quasiperiodic systems. It therefore provides general guidelines for realizing pure phases without MEs, as well as coexisting phases with MEs. For coexisting phases, one way to realize is to directly break the condition by adding higher QP frequencies to the system. Alternatively, one may break the chiral symmetry of the system by introducing spin-conserved hopping terms or QP chemical potential into the Hamiltonian footnote_chiral .
III.2 Universal mechanism for the emergence of critical states
We now show a new universal mechanism for the critical states in spinful QP chains: Spinful quasiperiodic systems can host critical states if the system possesses generalized incommensurate matrix element zeros ${\cal G}_{\Pi}$, which are defined as
${\cal G}_{\Pi}=\Big{\{}j_{k}\Big{|}\lim_{L\rightarrow\infty}p_{j_{k}}^{s}=0,\;s\in\{0,x,y,z\}\Big{\}}.$ (9)
This is a novel generalization of the mechanism for critical states in spinless QP chains to the spinful systems. The generalized incommensurate zeros (GIZs) in matrix elements refer to the incommensurately distributed zeros (IDZs) located in any component of the hopping coupling matrix $\Pi_{j}$ in the thermodynamic limit. The IDZs is defined as the site indices $\{j_{1},j_{2},\dots\}$ at which the hopping coefficients vanish in the thermodynamic limit ${\cal G}=\{j_{k}\mid\lim_{L\rightarrow\infty}t_{j_{k}}=0\}$ (Zhou et al., 2023; Simon and Spencer, 1989; Jitomirskaya and Marx, 2012), and the distribution of $\{j_{k}\}$ is incommensurate with the periodic of the system size. Similarly, the presence of GIZs in spinful models partitions the system into segments that remain invariant under dual transformations. This dual-invariance structure drives delocalized states into critical states [Fig. 2(b)].
The proof of this mechanism relies on the invariance of critical states under dual transformations. In the presence of the GIZs, the system includes processes $V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.}$, where $s$ and $s^{\prime}$ can be either identical or different. Under the dual transformation, these terms remain invariant as $V_{n}^{\mathrm{od}}c_{n+1s}^{\dagger}c_{ns^{\prime}}+\mathrm{h.c.}$ Consequently, when these terms dominate, critical orbitals enter the spectrum. From the perspective of the parent Hofstadter Hamiltonian, the spin-1/2 QP chain is mapped to a charged particle moving in a 2D bilayer system under a magnetic field. In this mapping, QP hopping in the hopping coupling matrix corresponds to diagonal hopping within or between the layers. This bilayer diagonal motion contributes dispersions along both $x$- and $y$-directions, thus introducing critical states into the projected 1D spectrum.
Inspired by this proof based on invariance under the dual transformation, we propose an additional guiding principle for generating critical states in spinful QP chains: introducing IDZs in the component of the on-site matrix shared with the hopping coupling matrix. For example, if the hopping coupling matrix are uniform, $\Pi_{j}\sim t\sigma_{+}$, then the system exhibits critical states if the on-site matrix $M_{j}$ exhibit IDZs in either the $\sigma_{x}$ or $\sigma_{y}$ component. This principle is demonstrated by the emergence and absence of critical states in mosaic lattice models and the new exactly solvable models we introduce in the next section. Notably, this principle is unique to spinful QP chains and does not extend to spinless QP chains. The key difference is that multiple generators in spinful QP chain allow for the transfer of IDZs from $M_{j}$ into either energy-dependent unbounded on-site potentials or IDZs in the hopping matrix, thereby generating critical states when reducing the spinful QP chain to a single chain. Some examples are provided in Theorem III in the next subsection.
The generalized incommensurate zeros (GIZs) ${\cal G}_{\Pi}$, and the incommensurately distributed zeros (IDZs) in the shared component of on-site and hopping coupling matrices serve as two guiding principles for introducing critical orbitals into the spectrum. These principles, together with the criteria for pure phases, provide powerful insights for constructing spinful QP models that host seven fundamental localization phases. For a generic spinful quasiperiodic system, Avila’s theory cannot analytically characterize the system. This motivates us to explore the conditions under which the spinful QP system becomes exactly solvable, as discussed in the following subsection.
III.3 Exact solvability of QP systems from local constraint
We now introduce the local constraint, with which the spinful QP system can exhibit exact solvability: A spinful quasiperiodic system can have exactly solvable points if the hopping coupling matrix $\Pi_{j}$ or its dual counterpart $\Pi_{n}$ are degenerate, which corresponds to
$\det|\Pi_{j}|=0\quad{\rm or}\quad\det|\Pi_{n}|=0,$ (10)
together with $p_{j}^{s}$ and off-diagonal terms of $M_{j}$ being either constant or purely quasiperiodic. This condition represents a local constraint, where the spinful quasiperiodic systems can be reduced to an effectively 1D spinless QP chain for the dressed particle with only nearest neighbor hopping, as illustrated in Fig. 2(c). In this case, Avila’s global theory can be applied to analytically characterize the localization properties of the dressed particles.
To elaborate the physics of this universal result, we consider the generic eigen-equation for the spinful QP system [Eq. (1)] $H\lvert\Psi\rangle=E\lvert\Psi\rangle$, with $\lvert\Psi\rangle=\sum_{j=1}^{L}(c^{\dagger}_{j,\uparrow}u_{j,\uparrow}+c^{\dagger}_{j,\downarrow}u_{j,\downarrow})\lvert\mathrm{vac}\rangle$, which is given by
$\Pi_{j-1}^{\dagger}\vec{u}_{j-1}+\Pi_{j}\vec{u}_{j+1}+M_{j}\vec{u}_{j}=E\vec{u}_{j}$ (11)
where $\vec{u}_{j}=(u_{j\uparrow},u_{j\downarrow})^{\intercal}$ denotes the spinnor for spin-1/2 particles. Under the constraint described in Eq.(10), the hopping coupling matrix takes one of the following fundamental forms: for spin-conserved processes, $\tilde{\Pi}_{j}\sim(\sigma_{0}\pm\sigma_{z})/2=\Lambda_{\pm}$; for spin-flipped processes, $\tilde{\Pi}_{j}\sim(\sigma_{x}\pm i\sigma_{y})/2=\sigma_{\pm}$; or a combination of both processes. Such constraint reduces the otherwise complicated eigen-equation in Eq.(11) to a 1D spinless QP chain with nearest-neighbor hopping. Here, $\tilde{\Pi}_{j}$ represents the hopping coupling matrix after a local unitary transformation, and the transformed spinnor is denoted by $\vec{v}_{j}$. We present the essential physics for the first two fundamental cases in the main text, and additional details for all three cases can be found in Appendix B.
For the spin-conserved constraint, $\tilde{\Pi}_{j}\sim t_{j}\Lambda_{\pm}$, the spin-up and spin-down particles can dress each other via
$v_{j,\downarrow}=\frac{M_{j}^{21}}{E-M_{j}^{22}}v_{j,\uparrow}.$ (12)
In this transformed basis, we denote $v_{j,\downarrow}=\psi_{j}$, and the original eigen-equation in Eq. (11) can be reduced to effective 1D chain with unchanged nearest-neighbor hopping, given by
$t_{j-1}\psi_{j-1}+t_{j}\psi_{j+1}+V_{j}^{\mathrm{eff}}\psi_{j}=E\psi_{j},$ (13)
where the effective energy-dependent on-site potential is
$V_{j}^{\mathrm{eff}}=M_{j}^{11}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}.$ (14)
Such on-site potential can serves as effective IDZs, which generates critical states in the system. By tailoring the on-site matrix, one can engineer an effective unbounded potential for the dressed particle, even when all couplings in the system are finite. This arises from resonant coupling between the system’s energy $E$ and the on-site modulation of the single species $M_{j}^{22}$, resulting in a divergent effective on-site potential. The divergence in the potential effectively partitions the 1D system into multiple sub-chains, manifesting IDZs in the hopping, and driving the delocalized eigenstates into critical states. Notably, this resonant coupling mechanism is generic, not limited to the exactly solvable regime, and can be extended to systems beyond the exactly solvable conditions.
For the spin-flipped constraint, $\tilde{\Pi}_{j}\sim t_{j}\sigma_{\pm}$, similarly, in the transformed basis
$v_{j,\downarrow}=\frac{t_{j}v_{j+1,\uparrow}+M_{j}^{21}v_{j,\uparrow}}{E-M_{j}^{22}},$ (15)
the original eigen-equation is reduced to
$t_{j-1}^{\mathrm{eff}}\psi_{j-1}+t_{j}^{\mathrm{eff}}\psi_{j+1}+V_{j}^{\mathrm{eff}}\psi_{j}=E\psi_{j},$ (16)
with effective hopping coefficient and on-site potential given by
$t_{j}^{\mathrm{eff}}=\frac{t_{j}M_{j}^{12}}{E-M_{j}^{22}},\quad V_{j}^{\mathrm{eff}}=M_{j}^{11}+\frac{t_{j-1}^{2}}{E-M_{j-1}^{22}}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}.$ (17)
These effective hopping and on-site potentials in Eqs. (14) and (17) highlight the requirement for $p_{j}^{s}$ and the off-diagonal elements of $M_{j}$ to be either constant or purely quasiperiodic in order to preserve exact solvability, as outlined in Theorem III (Sec. III.3). The effective eigen-equation for the dressed particles involves processes such as $M_{j}^{12}M_{j}^{21}$, $t_{j}M_{j}^{12}$, and $t_{j-1}^{2}$. These terms maintain a single QP frequency modulation when $p_{j}^{s}$ and off-diagonal elements of $M_{j}$ are either constant or purely quasiperiodic. However, when these terms are a mixture of constant and QP components, such as in the form $[A+B\cos(2\pi\alpha j)]^{2}$, they lead to mixed frequency modulations, thereby breaking the exact solvability.
The effective hopping coefficient and on-site potential in Eq. (17) explain the guiding principle from theorem II (Sec. III.2), where critical states arise if the system possesses IDZs in the shared components of both coupling matrices. Specifically, for the hopping coupling matrix $\Pi_{j}\sim t\sigma_{+}$, the presence of IDZs in $M_{j}^{12}$, which couples to $\sigma_{x}$ and/or $\sigma_{y}$, ensures that the effective 1D system exhibits IDZs, thus facilitating the emergence of critical states. This behavior is exemplified by the absence (or presence) of critical states in type-I (or type-II) mosaic lattice models Wang et al. (2020a); Zhou et al. (2023). For a comprehensive introduction to mosaic lattice models and their unification within the spin-1/2 QP framework, see Appendix C. The various types of mosaic lattice models can be unified as
$H_{\mathrm{M}}=\sum_{j}\lambda(c^{\dagger}_{j+1}\sigma_{-}c_{j}+\mathrm{h.c.})+\sum_{j}c_{j}^{\dagger}V_{j}^{\mathrm{M}}c_{j},$ (18)
where $c_{j}=(c_{j,\uparrow},c_{j,\downarrow})^{\intercal}$ is the spinor for the annihilation operators, $\lambda$ is the uniform hopping strengths. The onsite matrix $V_{j}^{\mathrm{M}}$ distinguishes the type-I and type-II mosaic lattices, respectively given by
$V_{j}^{\mathrm{M,I}}$ $=$ $2V_{0}V_{j}^{\mathrm{d}}\Lambda_{+}+\lambda\sigma_{x},$ (19)
$V_{j}^{\mathrm{M,II}}$ $=$ $2tV_{j}^{\mathrm{d}}(\sigma_{0}+\sigma_{x}),$ (20)
with $V_{0}$ and $t$ being the strength of QP modulation of the on-site matrix for type-I and type-II mosaic models, respectively. In the type-I mosaic lattice model, there are no IDZs in the shared components, resulting in a spectrum composed solely of extended and localized states Wang et al. (2020a). In contrast, for the type-II mosaic lattice model, IDZs appear on $\sigma_{x}$, which introduces energy-dependent quasiperiodic hopping and an unbounded on-site potential, thereby giving rise to rigorously defined critical states within the spectrum Zhou et al. (2023).
The universal results offer a powerful guidance for constructing exactly solvable models hosting various localization physics. The effective hopping and potential terms [Eq. (14)-(17)] provide a framework for designing the microscopic details of these systems by controlling the energy and coupling between the internal degrees of freedom. For example, by designing the effective potential, one can induce or suppress critical states, as well as manipulate the localization length.
Figure 3: New exactly solvable models constructed using universal results and dual transformations. The fractal dimension (FD) as a function of energy $E$ and hopping strengths $\lambda/t$. (a) The quasiperiodic spin-flipped (QPSF) model, derived by removing the $\sigma_{0}$ component of the type-II mosaic model, exhibits pure localized and critical phases. (b) The dual counterpart of the type-II mosaic lattice model, featuring analytic mobility edges (MEs) that separate extended and critical states. The MEs occur at $E_{c}=\pm\lambda$ and are marked by solid lines. (c) The dual counterpart of the QPSF model, which is equivalent to removing the $\sigma_{0}$ component in the dual model from (b). The dual-QPSF model exhibits pure extended and critical phases, with the phase transition points $\lambda=t$ marked by dashed lines. The system size is $L=2586$.
III.4 Examples of exactly solvable models based on universal results
With the guidance from the universal results, we introduce several new exactly solvable models derived from the type-II mosaic lattice model (Zhou et al., 2023) in Eq. (18) and Eq. (20) by removing the MEs. Through a combination of dual transformation, we can construct new nontrivial models, which demonstrate the application of the theorems. We investigate the phase diagram of the proposed model both analytically and numerically. We numerically solve the spectrum and employ the fractal dimension (FD) to phenomenologically characterize the localization properties of the eigenstates $\lvert\Psi\rangle=\sum_{j=1}^{L}u_{j}a_{j}^{\dagger}\lvert\mathrm{vac}\rangle$, where $\mathrm{FD}=-\lim_{L\rightarrow\infty}\ln\sum_{j=1}^{L}|u_{j}|^{4}/\ln L$, with $u_{j}$ being the wave function coefficients, $L$ the system size. In 1D, the FD approaches 1 for extended states and 0 for localized states, while for critical states, $10<\mathrm{FD}<$ 1. The FD quantifies the effective dimension experienced by an eigenstate: extended states uniformly spread across the system, yielding $\mathrm{FD}=1$. Localized states decay exponentially, effectively perceiving zero dimension, hence $\mathrm{FD}=0$. Critical states, however, exhibit self-similar wave function structures that span the entire system, resulting in an effective dimension between 0 and 1.
We begin with the type-II mosaic lattice model (Zhou et al., 2023) in Eq. (18) and Eq. (20), whose Hamiltonian is given by
$H_{\mathrm{M-II}}=\sum_{j}\lambda(c^{\dagger}_{j+1}\sigma_{-}c_{j}+\mathrm{h.c.})+2t\sum_{j}V^{\mathrm{d}}_{j}c_{j}^{\dagger}(\sigma_{0}+\sigma_{x})c_{j},$ (21)
where $\lambda$ is the uniform hopping strengths and $t$ is the strength of balanced QP potential and exchange coupling. The spectrum comprises critical (localized) states for the energies satisfy $E<|\lambda|$ ($E>|\lambda|$(Zhou et al., 2023).
Next, we construct new exactly solvable models that exhibit phase transitions between pure critical and localized phases by eliminating the MEs, as outlined in theorem I (Sec. III.1). This is accomplished by removing the spin-independent component of the on-site matrix. The resulting quasiperiodic spin-flipped (QPSF) model is given by
$H_{\mathrm{QPSF}}=\sum_{j}\lambda(c^{\dagger}_{j+1}\sigma_{-}c_{j}+\mathrm{h.c.})+2t\sum_{j}V_{j}^{\mathrm{d}}c_{j}^{\dagger}\sigma_{x}c_{j}.$ (22)
This model is exactly solvable, as the hopping coupling matrix satisfies $\det|\sigma_{\pm}|=0$. Applying Avila’s global theory, we find that the system is in the localized phase when QP spin-flipped process dominates $|t|>|\lambda|$, with the analytic localization length $\xi_{l}=2/\log|V/t|$ for all eigenstates. In contrast, when $|t|<|\lambda|$, the system is in the critical phase. The numerical results, as shown in Fig. 3(a), are consistent with the analytical results, with the dashed line marking the phase transition at $\lambda=t$.
This behavior can be understood as follows: In the limit $\lambda\rightarrow 0$, the QP on-site term localizes the system into dimmers with eigenenergies $E=\pm 2tV_{j}^{\rm d}$. In this case, the energy difference between adjacent site fluctuates significantly, and turning on the unifrom hopping $\lambda$ only hybridize a few dimmers with close energies, resulting in a localized phase. Conversely, in the limit $t\rightarrow 0$, the unifrom spin-flipped hopping creates degenerate dimers with flat band energies $E=\pm\lambda$. The degeneracy of these dimers is then lifted by the QP on-site term $V_{j}^{\rm d}$, which hybridizes them drastically, leading to a delocalized phase. The delocalized states are guaranteed to be critical due to the IDZs in the effective hopping
$t_{j}^{\mathrm{eff}}=\lambda V_{j}^{\mathrm{d}}/{E}.$ (23)
Furthermore, we develop new models by performing dual transformations to the two aforementioned models, which exhibit mobility edges or phase transitions between critical and extended states. Applying the dual transformation to the type-II mosaic model in Eq. (21) results in the following dual counterpart
$\mathrm{Dual}\big{[}H_{\mathrm{M-II}}\big{]}=\sum_{n}tc^{\dagger}_{n+1}(\sigma_{x}+\sigma_{0})c_{n}+\lambda c_{n}^{\dagger}V_{n}\sigma_{+}c_{n}+\mathrm{h.c.},$ (24)
with $V_{n}=\lambda\exp(i2\pi\alpha n)$. The model exhibits analytic MEs at $E_{c}=\pm\lambda$, with extended states for $E>|\lambda|$ and critical states for $E<|\lambda|$, as illustrated in Fig. 3(b). The absence of the localized orbitals can be understood as follows: under the local rotation $U=\exp(-i\pi\sigma_{y}/4)$, the coupling matrices of the dual model become
$\tilde{T}_{n}=2t\Lambda_{+},\quad\tilde{M}_{n}=\lambda\big{[}\cos(2\pi\alpha n)\sigma_{z}+\sin(2\pi\alpha n)\sigma_{y}\big{]}.$ (25)
In the limit $t\rightarrow 0$, the system yields two flat bands from the on-site matrix with energies $E=\pm\lambda$. The quantum states within these flat bands can be combined into either localized, extended, or critical states. The uniform hopping drives some states into extended states, while the IDZs in the shared component lead other states into critical states.
Similarly, the dual model of the QPSF model in Eq. (22) is given by
$\mathrm{Dual}\big{[}H_{\mathrm{QPSF}}\big{]}=\sum_{n}tc^{\dagger}_{n+1}\sigma_{x}c_{n}+\lambda c_{n}^{\dagger}V_{n}\sigma_{+}c_{n}+\mathrm{h.c.},$ (26)
with $V_{n}=\lambda\exp(i2\pi\alpha n)$, which is equivalent to removing spin-independent component from the previous dual model [Eq. (24)]. Through the duality transformation, this model exhibits a phase transition between pure critical and extended phases, with the critical phase for $|\lambda|>|t|$ and the extended phase for $|\lambda|<|t|$, with the correlation length given by $\xi_{c}=2/\log|V/t|$. The numerical results agree with the analytic predictions as shown in Fig. 3(c). This demonstrates Theorem III in Sec. III.3, which states that the model can be analytically characterized by investigating its dual counterpart.
We have demonstrated a direct application of the universal results to both explain existing models and develop new nontrivial ones. In the following section, we apply these theorems to construct novel models that incorporate all fundamental localization phases and MEs, providing a comprehensive phase diagram for the QP optical Raman lattice model.
IV Exactly solvable models for all fundamental localization phases
In this section, we further demonstrate the application of the universal results. By implementing these results, we construct new exactly solvable models with different types of mobility edges (MEs), including the QP optical Raman lattice model that hosts all seven fundamental phases. In particular, by combining the first two theorems (Sec.III.1 and Sec.III.2), we construct a model that hosts three pure phases and four mixed phases. We achieve this by utilizing the theorems to construct parameters that govern the existence of MEs and critical states. In the absence of critical states, we obtain two pure phases and their mixed phases: namely, pure extended, pure localized, and the mixture of extended and localized states, where the first two phases do not exhibit MEs, while the last one does. By introducing critical states, we obtain the remaining pure phase and three distinct mixed phases. Finally, we apply the third theorem in Sec. III.3 to derive the conditions for exact solvability of the QP optical Raman lattice model and to explain the full phase diagram of the model.
In the following, we construct the exactly solvable models hosting all fundamental MEs in Sec. IV.1 and the QP optical Raman lattice model hosting all seven fundamental phases in Sec. IV.2.
IV.1 The model hosting all fundamental MEs
We begin by constructing exactly solvable models hosting all fundamental MEs by manipulating the generalized incommensurate zeros (GIZs) in matrix elements and the incommensurately distributed zeros (IDZs) in the shared component. In the presence of the QP modulation on the hopping coupling, as shown in Fig. 4(a), we can anticipate the emergence of critical orbitals. In addition, the spin-independent component $\sigma_{0}$ in the GIZs or the shared component breaks chiral symmetry, indicating the existence of MEs. Thus, by combining critical states with broken chiral symmetry, we can construct a model that hosts all fundamental MEs. The unified Hamiltonian is given by
$\begin{split}H&=\sum_{j}\Big{[}c_{j+1}^{\dagger}\Big{(}t\Lambda_{+}+\mu V_{j}^{\mathrm{od}}\Lambda_{-}+\lambda_{1}\sigma_{x}\Big{)}c_{j}+\mathrm{h.c.}\Big{]}\\ &+\sum_{j}c_{j}^{\dagger}\Big{[}\big{(}V_{0}+2V_{B}V_{j}^{\mathrm{d}}\big{)}\Lambda_{-}+\lambda_{0}\sigma_{x}\Big{]}c_{j}.\end{split}$ (27)
This model shows how to manipulate the GIZs to introduce the critical orbitals into the system.
Figure 4: Generalized incommensurate zeros generated critical states. (a) The quasiperiodic spin-conserved hopping in this model gives rise to generalized incommensurate zeros (GIZs), which introduces critical orbitals in the system. (b) New exactly solvable model with generalized incommensurate zeros in matrix elements: Left panel: The model exhibits analytic mobility edges (MEs) at $E_{c}=\pm\lambda^{2}/\mu$, marked by the solid lines, which separate the extended states ($E<|\lambda^{2}/\mu|$) with fractal dimension ($\mathrm{FD}$) approaching 1, and the rigorous critical states ($E>|\lambda^{2}/\mu|$) with $\mathrm{FD}$ approaching a value between 0 and 1. Right panel: The corresponding $\mathrm{FD}$ for the dual model. The MEs $E_{c}=\pm\lambda^{2}/\mu$ (solid lines) now separate the localized states ($E<|\lambda^{2}/\mu|$) and the rigorous critical states ($E>|\lambda^{2}/\mu|$). (c) Exactly solvable model hosting all MEs: Upper panel: The energy-dependent effective potential resulting from the incommensurately distributed zeros in the shared component, which now acts as the IDZs on the effective dressed particle. Lower panel: The $\mathrm{FD}$ of the eigenstates shown as a function of $V_{B}/t$ and energy $E/t$, with $\lambda_{0}/t=1$ and $V_{0}/t=1.5$. The mobility edges are indicated by the solid lines. All systems have a size of $L=2586$.
We first introduce an exactly solvable model that hosts MEs separating critical and extended states, with the dual counterpart hosting MEs separating critical and localized states. By setting $t=V_{0}=V_{B}=0$ and $\lambda_{0}=\lambda_{1}=\lambda$, the hopping and on-site matrices become
$\Pi_{j}=\mu V_{j}^{\mathrm{od}}\Lambda_{-}+\lambda\sigma_{x},\quad M_{j}=\lambda\sigma_{x}.$ (28)
Fig. 4(b) left subfigure shows the model exhibits MEs separating critical and extended states at $E_{c}=\pm\lambda^{2}/\mu$. For energies $E<|\lambda^{2}/\mu|$, the system exhibits extended states, while for energies $E>|\lambda^{2}/\mu|$, critical states are observed. These results can be derived analytically by investigating the dual counterpart, whose coupling matrix is degenerate ($|\Pi_{n}|=0$) and is given by
$\begin{split}&\Pi_{n}=\mu V_{n}^{\mathrm{od}}\Lambda_{-},\\ &M_{n}=\lambda\big{[}\sigma_{x}+\cos(2\pi\alpha n)\sigma_{x}+\sin(2\pi\alpha n)\sigma_{y}\big{]}.\end{split}$ (29)
By reducing the system to an effective 1D spinless model, the effective hopping and on-site potential are given by
$t_{n}^{\mathrm{eff}}=\mu V_{n}^{\mathrm{od}},\quad V_{n}^{\mathrm{eff}}=\lambda^{2}(2+2V_{n}^{\mathrm{d}})/E.$ (30)
Neglecting the constant term in the potential, this effective model can be viewed as 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) without uniform hopping, where the system is in the localized (critical) phase when the QP on-site potential (QP hopping) dominates. Therefore, the system is in the critical phase for $\mu>\lambda^{2}/|E|$ and in the localized phase for $\mu<\lambda^{2}/|E|$. The numerical results shown in Fig. 4(b) right subfigure are consistent with this behavior. The MEs, $E_{c}=\pm\lambda^{2}/\mu$, separate the critical states for $|E|<\lambda^{2}/\mu$ from the localized states for $|E|>\lambda^{2}/\mu$. These results can also be derived by directly applying Avila’s global theory, yielding the Lyapunov exponent
$\gamma(E)=\frac{1}{2}\ln|\big{|}\lambda^{2}/\mu E\big{|}+\sqrt{(\lambda^{2}/\mu E)^{2}-1}\Big{|},$ (31)
which not only provides the MEs, but also gives the localization length $\xi_{l}=\gamma^{-1}(E)$ for the localized states.
These results in the dual space provide an analytic characterization in the original space. Specifically, for $E<|\lambda^{2}/\mu|$, the eigenstate with energy $E$ in real space is extended, with a correlation length $\xi_{e}=\gamma^{-1}(E)$, as localized states are transformed into extended states under the dual transformation. For $E>|\lambda^{2}/\mu|$, the eigenstate with energy $E$ remains critical.
We further introduce a new exactly solvable model which supports all three types of MEs. In addition to the MEs marking the critical-to-localized transition and critical-to-extended transition, the model also exhibits traditional MEs that separate the extended and localized states. By setting $\mu=\lambda_{1}=0$, the model becomes
$\Pi_{j}=t\Lambda_{+},\quad M_{j}=(V_{0}+2V_{B}V_{j}^{\mathrm{d}})\Lambda_{-}+\lambda_{0}\sigma_{x}.$ (32)
As guided by the principles, the incommensurately distributed zeros (IDZs) appear at the shared components $\sigma_{0}$ and $\sigma_{z}$, indicating the emergence of critical orbitals. As shown in Fig. 4(c), when $V_{B}/t$ is large enough, the spectrum consists of critical and localized states. The MEs $E_{c}=\pm 2t$ separate the critical states for $|E|<2t$ from the localized states for $|E|>2t$. For intermediate values of $V_{B}/t$, extended orbitals appear in the spectrum.
The model can be reduced to an effective 1D spinless model with the effective hopping and on-site potential given by
$t_{j}^{\mathrm{eff}}=t,\quad V_{j}^{\mathrm{eff}}=\lambda_{0}^{2}/(E-V_{0}-2V_{B}V_{j}^{\mathrm{d}}).$ (33)
This energy-dependent on-site potential, shown in the upper panel of Fig. 4(c), is key to understanding the phase diagram. When $|E-V_{0}|\leq 2V_{B}$, the system exhibits an energy-dependent unbounded potential, which acts as the IDZs, leading to only critical and localized orbitals. This explains why, when $V_{B}/t$ is sufficiently large, the system exhibits only critical and localized states. In this regime, the analytic Lyapunov exponent is given by
$\gamma(E)=\ln|\frac{|E/t|+\sqrt{(E/t)^{2}-4}}{2}\Big{|}.$ (34)
This directly provides the mobility edges, $E_{c}=\pm 2t$, which separate the critical states ($|E|<2t$) from the localized states ($|E|>2t$). This regime demonstrates the resonant coupling mechanism for critical states. Specifically, in the spinful QP chain, when the on-site flipped coupling is zero ($\lambda_{0}\to 0$), the spin-up and spin-down chains decouple. The spin-down chain produces the localized orbitals with energies $E_{j}=V_{0}+2V_{B}\cos(2\pi\alpha j)$, while the spin-up chain produces extended orbitals. and the spin-up chain produces the extended orbitals. Resonantly hybridizing these orbitals causes the effective on-site potential to diverge, leading to IDZs that generate critical states.
When $|E-V_{0}|>2V_{B}$, the effective on-site potential is finite, leading to extended and localized states in this regime. The LE is given by
$\gamma(E)=\max\Bigg{\{}\ln|\frac{\chi_{E}+\sqrt{\chi_{E}^{2}-\chi_{B}^{2}}}{|\chi_{0}|+\sqrt{\chi_{0}^{2}-\chi_{B}^{2}}}\Bigg{|},0\Bigg{\}},$ (35)
with $\chi_{B}=2V_{B}/t$, $\chi_{E}=EV_{B}/t^{2}$ and $\chi_{0}=(E-V_{0})/t$. A finite $\gamma(E)$ leads to localized states with localization length $\xi_{l}=\gamma^{-1}(E)$, while $\gamma(E)=0$ corresponds to extended states.
Finally, by combining transition condition for different effective potential and the corresponding LEs from Eq. (34) and Eq. (35), we determine the analytic MEs that separate extended, localized and critical states, as illustrated in Fig. 4(b) and summarized in Table 1.
Conditions Phases
$|E-V_{0}|>\max\{2V_{B},|E|V_{B}\}$ Extended
$2V_{B}>\max\{|E-V_{0}|,|E|V_{B}\}$ Critical
$|E|V_{B}>\max\{|E-V_{0}|,|2|V_{B}\}$ Localized
Table 1: The criteria for the eigenstates with energies $E$ to belong to one of the three phases: extended, critical, or localized, based on the relationship between the energy $E$, the on-site potential $V_{0}$, and the quasiperiodic modulation strength $V_{B}$.
IV.2 The model containing the seven fundamental localization phases
We further identify the seven fundamental phases realized in the 1D quasiperiodic optical Raman lattice model (Wang et al., 2020b, 2022b), as illustrated in Fig. 5(a). This model is constructed by manipulating the IDZs in the shared components and the chiral symmetry. Specifically, in the presence of chiral symmetry, three pure phases are obtained. Further breaking this symmetry introduces mobility edges and four additional coexisting phases. The Hamiltonian for the system is given by
$H$ $=\sum_{j}\Big{[}c_{j+1}^{\dagger}(t_{0}\sigma_{z}+it_{\mathrm{so}}\sigma_{y})c_{j}+\mathrm{h.c.}\Big{]}+$
$+M_{z}\sum_{j}c_{j}^{\dagger}\Big{[}\eta V_{j}^{\mathrm{d}}\sigma_{z}+(1-\eta)V_{j}^{\mathrm{d}}\sigma_{0}\Big{]}c_{j},$ (36)
where $t_{0}$ and $t_{\mathrm{so}}$ denote spin-conserved and spin-flipped hopping coefficients, $M_{z}$ is the strength of QP Zeeman potential, and $\eta$ is the chiral parameter that controls the ratio of spin-dependent to spin-independent QP potentials, governing the extent of chiral symmetry breaking.
Fig. 5(b) illustrates the seven fundamental phases by varying the QP Zeeman potential $M_{z}$ and the chiral parameter $\eta$. These results are obtained by numerically diagonalizing the QP optical Raman lattice model in Eq.(36) with $t_{\mathrm{so}}=0.8t_{0}$. We first examine two limiting cases: exact chiral symmetry ($\eta=1$) and complete chiral symmetry breaking ($\eta=0$), followed by a discussion of the intermediate value of $\eta$. In the case of complete chiral symmetry breaking $\eta=0$, corresponding to a purely spin-independent QP potential, the system exhibits extended and localized phases in the weak and strong quasiperiodic potential regimes, respectively. In this case, the transition energies remain outside the eigenstate spectrum, preserving the pure phases. For the intermediate value of $\eta$, a coexisting phase (L+E) emerges, with MEs separating extended and localized states. This behavior can be understood as follows: in the absence of spin-flipped hopping ($t_{\mathrm{so}}=0$), the model in Eq.(36) describes two decoupled spin-up and spin-down QP chains with opposite energy spectra, where each chain hosts either purely localized or extended phases. Turning on $t_{\rm so}$ couples the two chains, opening gaps within each and flattening the dispersion of original band states. This coupling localizes part of states in the moderate potential regime, leading to the formation of MEs. On the other hand, in the chiral symmetry limit $\eta=1$, which corresponds to purely QP Zeeman potential, distinct phases appear in the weak, moderate, and strong field regimes, corresponding to pure extended, critical, and localized phases, respectively. These results are consistent with the predictions of Theorem I (Sec.III.1) and previous studies Wang et al. (2016, 2020b). As indicated by theorem I and II, coexisting phases involving critical states (C+E, L+C, L+E+C) necessitate introducing generalized incommensurate zeros in matrix elements and breaking chiral symmetry. This corresponds to the presence of both spin-dependent and spin-independent QP potentials. Fig. 5(b) indicates that these phases occur for $\eta_{c}<\eta<1$, with $\eta_{c}\approx 0.43$.
The emergence of the seven fundamental phases can be analytically predicted by combining the universal results and Avila’s global theory. By applying Theorem III in Sec. III.3, the system is reduced to a spinless 1D QP chain with nearest neighbor hopping when $\det|\Pi_{j}|=0$, which occurs when $|t_{0}|=|t_{\mathrm{so}}|$. We set $t_{0}=t_{\mathrm{so}}$ without losing generality. The system exhibits chiral symmetry at $\eta=1$, where only pure critical and localized phases exist when $t_{\rm so}=t_{0}$, and all three pure phases are present when $t_{\rm so}\neq t_{0}$. Introducing a spin-independent quasiperiodic potential ($\eta\neq 1$) breaks chiral symmetry, leading to another exactly solvable point at $\eta=1/2$, where MEs emerge within the original critical and localized phases for $t_{\rm so}=t_{0}$, giving rise to the L+C phase. Further, various coexisting phases with combinations of the three types of quantum states emerge for $\eta<1$ when $t_{\mathrm{so}}\neq t_{0}$. Therefore, both numerical and analytical results confirm the existence of the seven fundamental localization phases, establishing this system as a universal quantum platform for exploration of localization physics.
In the following, we elaborate the entire phase diagram, starting from the high symmetry line $t_{0}=t_{\mathrm{so}}$, where the effective nearest-neighbor hopping is given by
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\eta\Delta_{j}}{E-(1-\eta\Delta_{j})},$ (37)
with $\Delta_{j}=M_{z}V_{j}^{\mathrm{d}}$. The energy-dependent effective on-site potential is given by
$V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}}{E-(1-\eta)\Delta_{j-1}}+\frac{E(1-\eta)\Delta_{j}-(1-2\eta)\Delta_{j}^{2}}{E-(1-\eta)\Delta_{j}}.$ (38)
We first analyze the three limiting cases, $\eta=0$, $\eta=1$, and $\eta=1/2$ while keeping $t_{0}=t_{\rm so}$, and then extend the discussion to cases where $\eta$ deviates from these limits and $t_{0}\neq t_{\mathrm{so}}$, to explain the entire phase diagram.
Figure 5: Seven fundamental phases of localization physics realized in 1D quasiperiodic (QP) optical Raman lattice. (a) Model Illustration: The hopping coupling matrix includes both spin-conserved and spin-flipped hopping terms, while the on-site matrix incorporates both spin-dependent quasiperiodic (QP) potential coupled to $\sigma_{z}$ and spin-independent QP potential coupled to $\sigma_{0}$. (b) Phase Diagram: The phase diagram, shown as a function of the QP Zeeman potential $M_{z}$ and the chiral parameter $\eta$, reveals seven distinct phases. These consist of three pure phases: extended (E), localized (L), and critical (C), as indicated in the diagram. In addition, there are three coexisting two-phase regions: (L+E) for the coexistence of localized and extended states, (L+C) for localized and critical states, and (C+E) for critical and extended states. Moreover, a coexisting three-phase region (L+E+C) appears, where all three phases coexist. The phase diagram is obtained with the parameter set $t_{\mathrm{so}}=0.8t_{0}$. (c) Exact Solvable Points: (c1) When $\eta=0$, the entire spectrum is localized. (c2) For $\eta=0.5$, the spectrum splits into localized and critical states, with the mobility edges marked by solid lines at $E_{c}=\pm 2t_{0}$. States with $|E|<2t_{0}$ are critical, while those with $|E|>2t_{0}$ are localized. (c3) At $\eta=1$, the system exhibits pure localized and critical phases, with the transition point marked by the dashed line at $M_{z}=4t_{0}$. The system is in the localized (critical) phase when $M_{z}>4t_{0}$ ($M_{z}<4t_{0}$). The system size used in these calculations is $L=2586$.
IV.2.1 Pure spin-independent quasiperiodic potential
For pure spin-independent limit of QP potential, where $\eta=0$, the system exhibits only extended and localized states, with no critical states since IDZs in the shared component vanish. We first consider the high-symmetry case where $t_{0}=t_{\mathrm{so}}$, which leads to a vanishing effective hopping, leaving only the on-site potential
$t_{j}^{\mathrm{eff}}=0,\quad V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}}{E-\Delta_{j-1}}+\Delta_{j}.$ (39)
In this case, all eigenstates are localized as shown in Fig. 5(c1). The mechanism behind this can be understood as follows: when $M_{z}=0$ (i.e., no QP Zeeman potential), the Hamiltonian [Eq. (36)] yields two flat bands, and any finite QP onsite potential $M_{z}$ can fully localize all states. Deviating from the high-symmetry line $t_{0}\neq t_{\mathrm{so}}$ introduces next-nearest-neighbor (NNN) hopping, which is given by
$t_{j}^{\mathrm{NN}}=\frac{t_{+}t_{-}}{E-\Delta_{j+1}},$ (40)
where $t_{\pm}=t_{0}\pm t_{\mathrm{so}}$. The nearest-neighbor hopping remains zero, and on-site potential is modified to
$t_{j}^{\mathrm{eff}}=0,\quad V_{j}^{\mathrm{eff}}=\frac{t_{+}^{2}}{E-\Delta_{j-1}}+\frac{t_{-}^{2}}{E-\Delta_{j+1}}+\Delta_{j}.$
The emergence of this NNN hopping begins to destabilize localization, allowing for the appearance of delocalized states. Although the system is no longer analytically solvable for $t_{0}\neq t_{\rm so}$, the possible phases can still be determined. For sufficiently large (weak) $M_{z}$, the system always enters the localized (extended) phase. On the other hand, for intermediate $M_{z}$ values, the chiral symmetry is explicitly broken by spin-independent QP modulation $M_{z}V_{j}^{\mathrm{d}}\sigma_{0}$, leading to the emergence of mobility edges (MEs) between the extended and localized states. This accounts for extended, localized and L+E phases in Fig. 5(b) at $\eta=0$, which extend to small $\eta$ regime with $\eta<\eta_{c}\approx 0.43$.
IV.2.2 Pure spin-dependent quasiperiodic potential
For pure spin-dependent QP potential, where $\eta=1$, we begin by considering the scenario where $t_{0}=t_{\mathrm{so}}$, which yields the following effective hopping and on-site potential
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\Delta_{j}}{E},\quad V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}+\Delta_{j}^{2}}{E}.$ (41)
The corresponding Lyapunov exponent (LE) is given by
$\gamma=\max\Big{\{}\frac{1}{2}\ln|\frac{M_{z}}{4t_{0}}\Big{|},0\Big{\}}.$ (42)
When $|M_{z}|>4|t_{0}|$, the QP potential dominates, leading to a localized phase, where all states exhibit a localization length $\xi=\gamma^{-1}$. When $|M_{z}|<4|t_{0}|$, the LE vanishes ($\gamma=0$), and the system enters the critical phase. This behavior is attributed to the IDZs in the hopping coefficients, consistent with the numerical calculations shown in Fig. 5(c3). No extended states are observed in this regime when $\eta=1$ and $t_{0}=t_{\mathrm{so}}$ as expected.
We now extend this analysis to the case where $t_{0}\neq t_{\mathrm{so}}$, which modifies the effective nearest-neighbor hopping and on-site potential due to the imbalance between $t_{0}$ and $t_{\mathrm{so}}$
$t_{j}^{\mathrm{eff}}=\frac{t_{-}\Delta_{j+1}-t_{+}\Delta_{j}}{E},\quad V_{j}^{\mathrm{eff}}=\frac{2t_{0}^{2}+2t_{\mathrm{so}}^{2}+\Delta_{j}^{2}}{E},$
and more importantly, the imbalance introduces an effective next-nearest-neighbor hopping
$t_{j}^{\mathrm{NN}}=t_{+}t_{-}/E.$ (43)
This NNN hopping, induced by the imbalance between $t_{0}$ and $t_{\mathrm{so}}$, gives rise to extended orbitals within the phase diagram. When the imbalance $|t_{0}-t_{\mathrm{so}}|$ is small relative to $M_{z}$, the critical states, generated by zeros in the hopping terms, remain dressed by the NNN hopping $t_{j}^{\mathrm{NN}}$, preserving the critical phase. However, as the imbalance increases and $t_{j}^{\mathrm{NN}}$ dominates, the critical states transition into extended states Huang et al. (2025).
The system exhibits the chiral symmetry at $\eta=1$, implying the absence of MEs and the presence of only three pure phases: localized, extended, and critical. The global theory does not strictly apply here, however, because the system has no mobility edges, the phase boundaries can be determined by analyzing the typical eigenstates. For example, by considering the zero-energy states $E=0$, the eigen-equation Eq. (36) becomes
$-t_{0}(\psi_{j+1}+\psi_{j-1})+t_{\mathrm{so}}(-\psi_{j+1}+\psi_{j-1})+\Delta_{j}\psi_{j}=0,$
$-t_{0}(\varphi_{j+1}+\varphi_{j-1})+t_{\mathrm{so}}(\varphi_{j+1}-\varphi_{j-1})+\Delta_{j}\varphi_{j}=0.$
Here $\psi_{j}$ and $\varphi_{j}$ are the wavefunction of the spin-up and spin-down states, and the eigen-equation is expressed as the form of transfer matrix $\big{(}\begin{array}[]{cc}\psi_{j+1},&\psi_{j}\end{array}\big{)}^{\intercal}=A_{j}\big{(}\begin{array}[]{cc}\psi_{j},&\psi_{j-1}\end{array}\big{)}^{\intercal}$ and $\big{(}\begin{array}[]{cc}\varphi_{j+1},&\varphi_{j}\end{array}\big{)}^{\intercal}=B_{j}\big{(}\begin{array}[]{cc}\varphi_{j},&\varphi_{j-1}\end{array}\big{)}^{\intercal}$ where the transfer matrix $A_{j}$ is
$A_{j}=\left(\begin{array}[]{cc}\Delta_{j}/{t_{+}}&-t_{-}/t_{+}\\ 1&0\end{array}\right),$
and $B_{j}$ has a similar expression. In the spin-dependent limit ($\eta=1$), the delocalized phase is topological while the localized phase is topologically trivial Wang et al. (2020b). The phase boundary between the critical and localized phases can thus be determined by the topological transition. If both eigenvalues of the transfer matrix $A=\prod_{j=1}^{L}A_{j}$ are either less than 1 or greater than 1, the system is topological, and the $\psi$ zero mode is localized at one end of the chain. Assuming $t_{0}>0$ and $t_{\mathrm{so}}>0$, then the two eigenvalues of $A$ satisfy $|\lambda_{1}\lambda_{2}|<1$ , the topological nature is dictated by the larger eigenvalue $|\lambda_{2}|$. After performing a standard similarity transformation DeGottardi et al. (2011, 2013) or applying the global theory, we find the localized-to-critical transition point is given by $|M_{z}/2t_{+}|=1$. By applying the same reasoning to the transfer matrix in the dual space, the phase boundary between localized and critical phases in the dual space is $|M_{z}/2t_{-}|=1$, corresponding to the transition between extended and critical phases in the original space. Therefore, for $\eta=1$ and $t_{0}\neq t_{\mathrm{so}}$, the system exhibits three distinct phases as summarized in Table. 2.
Conditions Phases
$M_{z}<2|t_{-}|$ Extended
$2|t_{-}|<M_{z}<2|t_{+}|$ Critical
$M_{z}>2|t_{+}|$ Localized
Table 2: The criteria for the phases of 1D QP optical Raman lattice model at $\eta=1$, based on the relationship between the QP Zeeman potential $M_{z}$ and the imbalanced hopping $t_{\pm}=t_{0}\pm t_{\mathrm{so}}$.
When chiral parameter $\eta$ deviates from $\eta=1$, the chiral symmetry is broken and MEs appear. This leads to the emergence of the C+E phase, which interpolates between the pure critical and extended phases, and the L+C phase, which interpolates between the pure localized and critical phases, as shown in Fig. 5(b). Furthermore, as the chiral parameter $\eta$ is further deviated from 1, the L+E+C phase emerges, interpolating between the C+E and L+E phases.
IV.2.3 Balanced quasiperiodic potential
For balanced QP potential with $\eta=1/2$, the energy-dependent effective nearest-neighbor hopping and on-site potential are given by
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\Delta_{j-1}}{2E-\Delta_{j-1}},\quad V_{j}^{\mathrm{eff}}=\frac{8t_{0}^{2}}{2E-\Delta_{j-1}}+\frac{E\Delta_{j}}{2E-\Delta_{j}}.$ (44)
Applying Avila’s theory, the LE for the system is
$\gamma(E)=\max\Big{\{}\frac{1}{2}\ln||E/2t_{0}|+\sqrt{E^{2}/4t_{0}^{2}-1}\Big{|},0\Big{\}}.$ (45)
The states with energy $|E|>2|t_{0}|$ are localized, with localization length $\xi(E)=\gamma^{-1}(E)$. The states are critical when $|E|<2|t_{0}|$, as the $\gamma(E)=0$ and the hopping terms have IDZs. Consequently, $E=\pm 2t_{0}$ are critical energies separating localized and critical states, indicating the presence of MEs as shown in Fig. 5(c2). Thus, for $\eta=1/2$ and $t_{0}=t_{\mathrm{so}}$, the system is in L+C phase. We extend this analysis to the case where $t_{0}\neq t_{\mathrm{so}}$, introducing an effective NNN hopping term given by
$t_{j}^{\mathrm{NN}}=\frac{-2t_{+}t_{-}}{2E-\Delta_{j+1}}.$ (46)
Similarly, the imbalance between $t_{0}$ and $t_{\mathrm{so}}$ generates extended orbitals within the spectrum, while critical states still exist when the NNN hopping $t_{j}^{\mathrm{NN}}$ is small relative to $M_{z}$. The combination of extended orbitals and the L+C phase results in pure extended, L+E, and L+E+C phases. Similar phase transitions also occur when $\eta$ deviates from $\eta=1/2$. This analysis clarifies the origin of the L+C phase at the exactly solvable point $\eta=1/2$ and $t_{0}=t_{\mathrm{so}}$, and demonstrates how the related phases,extended, critical, L+E, and L+E+C phases, emerge near $\eta=1/2$, as depicted in Fig. 5(b).
Figure 6: Experimental realization schemes. (a)-(b) Implementation of the model supporting all fundamental mobility edges (MEs) (a) The spin-conserved hopping for spin-down atoms is suppressed by a spin-dependent deep primary potential ${\cal V}_{p}(z)$ (blue solid line), while spin-up atoms undergo hopping in a secondary potential ${\cal V}_{s}(z)$ (red solid line). The on-site potential for spin-down atoms (red dashed line) is realized by the combination of ${\cal V}_{p}(z)$ and ${\cal V}_{s}(z)$ , which are incommensurate with each other. The on-site spin-flipped process is introduced via an overall coupling $M_{0}$. (b) Schematic of the experimental setup. The spin-dependent potential ${\cal V}_{p}(z)$ is formed by the combination of ${\bf E}_{2}$ and ${\bf E}_{3}$ with the traveling wave ${\bf E}_{4}$, while the spin-independent secondary potential ${\cal V}_{s}(z)$ is generated by ${\bf E}_{1}$, affecting both spin states. (c)-(d) Realization of the quasiperiodic optical Raman lattice model. (c) The spin-conserved hopping $t_{0}$ is induced by the primary lattice ${\cal V}_{p}(z)$, and the spin-flipped hopping $t_{\rm so}$ is induced by the Raman potential ${\cal M}(z)$. The quasiperiodic Zeeman potential is realized by the secondary lattice ${\cal V}_{s}(z)$. (d) Schematic of the setup. A standing wave ${\bf E}_{1}$, combined with the blue-detuned spin-independent primary lattice generated by ${\bf E}_{2}$, creates the spin-dependent Zeeman potential. A traveling wave ${\bf E}_{3}$ generates the Raman potential ${\cal M}(z)$ in combination with ${\bf E}_{2}$.
IV.3 Unified picture for the seven fundamental localization phases
It is important to highlight the mapping relation between two analytically solvable models: the one used in our experiment and the type-II mosaic lattice model Zhou et al. (2023) defined in Eq. (21), with slightly modified on-site potential, given by
$H_{\mathrm{M}}=\sum_{j}\lambda(c^{\dagger}_{j+1}\sigma_{-}c_{j}+\mathrm{h.c.})+\sum_{j}V^{\mathrm{d}}_{j}c_{j}^{\dagger}(2V_{0}\sigma_{0}+2t\sigma_{x})c_{j}.$ (47)
Here, $V_{0}$ is the strength of the on-site potential. At the high symmetry line $V_{0}=t$, the model Eq. (47) has analytic expression for the MEs as discussed in Sec. III.4. Its MEs between critical and localized states, and this mapping provides a systematic framework for understanding the entire phase diagram.
The QP optical lattice model model at $t_{0}=t_{\mathrm{so}}$ can be mapped to the type-II mosaic lattice model [Eq. (47)] followed by a local rotation $U=\exp(-i\pi\sigma_{y}/4)$. The coefficients between the two models are related by the following substitution rule
$\lambda$$\leftrightarrow$ $2t_{0},$ (48)
$2t$$\leftrightarrow$ $\eta M_{z},$ (49)
$2V_{0}$$\leftrightarrow$ $(1-\eta)M_{z}.$ (50)
In the case of a pure spin-dependent QP potential ($\eta=1$) of QP optical Raman lattice model, the type-II mosaic model reduces to the uniform on-site potential limit $V_{0}=0$. In this scenario, both models are exactly solvable, exhibiting localized and critical phases. Specifically, the type-II mosaic model is in the localized phase when $|t|>|\lambda|$ and in the critical phase when $|t|<|\lambda|$ [Fig. 3(a)]. This is consistent with the QP optical Raman lattice model, where the localized phase occurs when $M_{z}>4t_{0}$ and the critical phase when $M_{z}<4t_{0}$ [Fig. 5(c3)].
For the balanced quasiperiodic potential ($\eta=1/2$), the type-II mosaic model is along the high-symmetry line $V_{0}=t$. In this case, both models are exactly solvable again and exhibits only L+C phase, which corresponds to the coexistence of localized and critical states due to the presence of mobility edges. The type-II mosaic model exhibits exact MEs at $E=\pm\lambda$ separating the localized and critical states Zhou et al. (2023), which corresponds to the MEs which are given by $E=\pm 2t_{0}$ in the QP optical Raman lattice model.
Reducing the chiral parameter $\eta$ from 1 to 1/2 corresponds to turning on the on-site potential $V_{0}$ in the type-II mosaic model, transitioning from $V_{0}=0$ to $V_{0}=t$. Breaking the chiral symmetry of both models, introducing the MEs between localized and critical states, leading to the emergence of L+C phase. As the on-site potential increases towards the high-symmetry line $V_{0}=t$, or when the QP optical Raman lattice model reaches $\eta=1/2$, the system becomes analytically solvable again, exhibiting L+C phase in the phase diagram and indicating the presence of MEs between localized and critical states.
Further introducing the imbalance between $t_{0}$ and $t_{\mathrm{so}}$ leads to the emergence of extended phases and associated MEs. And this will lead to the long-range hopping terms when reduced the system into the spinless QP chain, introducing extended orbital into the spectrum of the type-II mosaic model. Thus, this mapping, combined with the imbalance between $t_{0}$ and $t_{\mathrm{so}}$, accounts for the generation of the seven phases in the phase diagram.
V Schemes for experimental realization
In this section, we propose experimental schemes for realizing models that host all three fundamental types of mobility edges (MEs) [Eq.(32)] and all seven fundamental phases [Eq.(36)]. These schemes utilize ultracold alkali atoms, which are ideal candidates due to their large fine-structure energy splitting and moderate natural linewidth. These properties facilitate efficient Raman coupling and spin-dependent potentials, even when the laser frequency is far detuned from the $D_{1}$ and $D_{2}$ lines. Such features have been extensively studied both theoretically and experimentally Soltan-Panahi et al. (2011); Kuzmenko et al. (2019); Szulim et al. (2022), making ultracold alkali atoms well-suited for implementing an incommensurate lattice schemes.
We first propose the scheme for the model in Eq. (27) with $\mu=\lambda_{1}=0$ supporting all three fundamental types of MEs, as illustrated in Fig. 6(a-b), whose Hamiltonian is given by
$H=\left[\frac{p_{z}^{2}}{2m}+{\cal V}_{s}(z)\right]\otimes\sigma_{0}-{\cal V}_{p}(z)\Lambda_{-}+M_{0}\sigma_{x},$ (51)
where the first term represents the kinetic energy. Here ${\cal V}_{p}(z)$ is a deep spin-dependent primary lattice that freezes the spin-conserved hopping for spin-down atoms only, and is given by ${\cal V}_{p}(z)=V_{p}\cos(k_{p}z+\phi_{p})$. The terms in Eq. (32) can be realized as follows. The uniform spin-conserved hopping term $\Pi_{j}=t\Lambda_{+}$ for spin-up atoms can be obtained by
[…]
where $w_{s,\uparrow}(z)$ is the Wannier function of the secondary lattice generated by spin-independent potential ${\cal V}_{s}(z)=V_{s}\cos(k_{s}z+\phi_{s})$, with $a_{s}=\pi/k_{s}$ being the lattice constant. The potential ${\cal V}_{s}(z)$ is incommensurate with the potential ${\cal V}_{p}(z)$ for spin-down atoms, leading to the QP potential $M_{j}=2V_{B}V_{j}^{\rm d}\Lambda_{-}$ for spin-down atoms
[…]
The Hamiltonian in Eq. (51) can be realized based on QP optical Raman lattice, as outlined below. The detailed implementation can be found in Appendix. E. To facilitate the description, we perform a unitary transformation, such that $\sigma_{z}\rightarrow\sigma_{x}$ and $\sigma_{x}\rightarrow-\sigma_{z}$. The kinetic term and spin-independent potential ${\cal V}_{s}(z)$ remain unchanged and can be generated using the standing wave ${\bf E}_{1}$. The rotated spin-dependent potential ${\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2$ can be realized using a standing waves ${\bf E}_{2}$ and ${\bf E}_{3}$ along with a traveling wave ${\bf E}_{4}$. Finally, the $M_{0}\sigma_{z}$ term arises from the two-photon detuning of the Raman coupling. After performing the inverse unitary trnasformation, we reach the Hamiltonian in Eq. (51). More details can be found in Appendix. E, where ${}^{87}{\rm Rb}$ atoms are used as an exmaple.
We then propose the scheme for the model supporting all seven fundamental phases [Eq. (36)] based on the QP optical Raman lattice [Fig. 6(c,d)], which has been successfully employed for the realization of synthetic gauge fields and the exploration of topological phases with ultracold atoms Zhang and Liu (2018); Liu et al. (2014); Wu et al. (2016); Wang et al. (2018); Sun et al. (2018); Song et al. (2019); Lu et al. (2020); Wang et al. (2021c); Zhang et al. (2024). The Hamiltonian is given by
$\begin{split}H&=\bigg{[}\frac{p_{z}^{2}}{2m}+{\cal V}_{p}(z)+(1-\eta){\cal V}_{s}(z)\bigg{]}\otimes\sigma_{0}\\ &+{\cal M}(z)\sigma_{x}+\eta{\cal V}_{s}(z)\sigma_{z}.\end{split}$ (54)
The different terms in Eq. (36) can be realized accordingly. The uniform hopping coupling $\Pi_{j}=t_{0}\sigma_{z}+it_{\rm so}\sigma_{y}$ arises from the tunneling of the Wannier function of the primary lattice, where
[…]
with $p_{z}^{2}/2m$ being the kinetic energy along the lattice direction $z$, $a_{0}$ the primary lattice constant, and ${\cal V}_{p}(z)=V_{p}\cos^{2}(k_{p}z)$ the spin-independent primary lattice realized by a blue-detuned standing wave ${\bf E}_{2}$. The $t_{\rm so}$ is generated by the Raman potential
[…]
where ${\cal M}(z)=M_{0}\cos(k_{p}z)$ represents the Raman potential formed by the intersection of $x$-polarized ${\bf E}_{2}$ and $z$-polarized traveling wave ${\bf E}_{3}$, generating the spin-flipped hopping. The QP on-site matrix element is realized by the spin-dependent secondary lattice, which is given by
[…]
where the QP parameter $\alpha=k_{s}/k_{p}$, and ${\cal V}_{s}(z)=V_{s}\cos^{2}(k_{s}z+\phi_{s})$ is the spin-dependent secondary lattice generated by the composite standing wave ${\bf E}_{1}$, formed by two counter-propagating beams with mutually perpendicular polarization along the $x$ and $y$ axes. The chiral parameter $\eta$ is tuned by adjusting the power ratio of the two different polarizations of beam ${\bf E}_{1}$. The spin-dependent term $\eta{\cal V}_{s}(z)\sigma_{z}$ from vector optical shift is induced by the y-polarization, while the spin-independent term $(1-\eta){\cal V}_{s}(z)\sigma_{0}$ is mainly contributed from the scalar components of both the x- and y-polarizations.
VI Conclusion and outlook
We have proposed a generic spin-$1/2$ quasiperiodic (QP) system that unifies the existing important 1D QP models, some of which requiring an extra Majorana representation, and established a unified framework for all fundamental localization phases in quasiperiodic systems. The framework is built on three universal results. First, we showed with renormalization group method the criteria for obtaining the pure phases, namely when the system has no mobility edges. Second, we uncovered a new universal mechanism for the emergence of critical states, as characterized by the generalized incommensurate zeros in matrix elements that remain invariant under the dual transformation and introduce critical orbitals into the system. Third, we identified the condition for the exact solvability of QP systems from local constraint, which also provides the mechanism to realize effective unbounded on-site potential in spinful QP chains. These results establish a unified theory which enables an in-depth exploration of the broad classes of all fundamental localization phases in quasiperiodic systems, and offer the key insights for constructing their exactly solvable models. In particular, we proposed a novel model which hosts all basic types of mobility edges (MEs), and further predict a highly nontrivial quasiperiodic system which hosts all the seven fundamental phases in Anderson localization physics. Finally, we proposed experimentally feasible schemes to realize the theoretical predictions, both in Rydberg atom systems and in optical Raman lattice.
This work enables a broad exploration and exact characterization of the fundamental quantum phases in Anderson localization and opens up intriguing topics for the future study. For example, while the present study is focused on the spin-$1/2$ QP systems, it is natural to extend the current study to the QP systems with larger spins, e.g. to the 1D $\mathrm{SU}(N)$ QP chain with $N$ larger than $2$. Generalizing the mechanism of IDZs in matrix elements for critical states to the $\mathrm{SU}(N)$ QP systems is of great interests to obtain the rigorous critical states in generic QP systems. The exact characterization of the distinct localization phases in such $\mathrm{SU}(N)$ QP system could potentially be connected to the exactly solvable mosaic lattice models (Wang et al., 2020a; Zhou et al., 2023; Longhi, 2024a) with the large unit cell, and deserves particular future efforts. On the other hand, extending the present results to higher dimensions is another important issue for the future study.
Moreover, the newly proposed exactly solvable models in this work, which host various MEs, can further be implemented in the many-body regime. The different types of the MEs in this work, including the traditional ME separating localized and extended states, and the novel MEs involving critical states, can be considered to study the interplay between many-body localization, ergodic and many-body critical phases. The exact solvability of the model, which allows for the analytic characterization of the whole non-interacting spectrum, serves as an analytic platform to study the interplays among them.
Acknowledgements.
This work was supported by National Key Research and Development Program of China (Grants No. 2021YFA1400900, No. 2022YFA1405800 and No. 2020YFA0713300), the National Natural Science Foundation of China (Grants No. 12425401, No. 12261160368 and No. 12401208), the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000), the Shanghai Municipal Science and Technology Major Project (Grant No. 2019SHZDZX01), the Natural Science Foundation of Jiangsu Province (Grant No. BK20241431), the Nankai Zhide Foundation, and by the Fundamental Research Funds for the Central Universities (Grant No. 100-63253099).
Appendix A Proof of criteria for the pure phases without MEs
In this section, we provide the detail of the proof of the criteria for the system exhibits pure phase and exhibits no mobility edges (MEs), as discussed in the main text, using the renormalization group technique.
Under the commensurate approximation of the irrational parameter, the system displays periodic structure and exhibit band dispersion. And the characteristic polynomial is given by the determinant $P^{(n)}(E;\kappa_{x},\kappa_{y})=|{\cal H}^{(n)}-E|$, which can be expanded over the principle dispersion
$P^{(n)}(E;\kappa_{x},\kappa_{y})$ $=t_{R}^{(n)}\cos(\kappa_{x}+\kappa_{x}^{0})+V_{R}^{(n)}\cos(\kappa_{y}+\kappa_{y}^{0})$
$+\mu_{R}^{(n)}\cos(\kappa_{x}+\tilde{\kappa}_{x}^{0})\cos(\kappa_{y}+\tilde{\kappa}_{y}^{0})$
$+\epsilon_{R}^{(n)}(E,\varphi,\kappa)+T_{R}^{(n)}(E).$ (58)
The system exhibits pure phases without MEs correspond to the case that, the coefficients associated with the relevant dispersions are energy independent. In the presence of the commensurate approximation, then the band dispersion is periodic with respect to both $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. Here, $k_{x}$ is the twisted momentum attached to the hopping coupling matrix $\Pi_{j}\rightarrow\Pi_{j}e^{ik_{x}}$ and $k_{y}$ is the phase offset in the QP modulation $V_{j}^{\mathrm{d}}$ and $V_{j}^{\mathrm{od}}$. Since $P(E)$ has only real roots as $H$ is Hermitian, so the dispersion and $P(E;\kappa_{x},\kappa_{y})$ shares similar periodicity, thus $P(E;\varphi,\kappa)$ is also periodic with respect to $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$.
As indicated in the main text, we first diagonalize the on-site matrix $M_{j}$ as $M_{j}=m_{j}^{z}\sigma_{z}+m_{j}^{0}\sigma_{0}$, and the resulted hopping coupling matrix is still in the form $\Pi_{j}=\sum_{s}p_{j}^{s}\sigma_{s}$, with $s$ being $s=\{0,x,y,z\}$. For the criteria that the coefficients of hopping coupling matrix satisfy
$p_{j}^{x}/p_{j}^{y}=\mathrm{Const.}\in\mathbb{R},$ (59)
and both coupling matrices exclude $\sigma_{0}$, namely
$\Pi_{j}=p_{j}^{x}\sigma_{x}+p_{j}^{y}\sigma_{y}+p_{j}^{z}\sigma_{z},\quad M_{j}=m_{j}^{z}\sigma_{z}.$ (60)
We can first simplify model by denoting $m_{j}^{z}\equiv m_{j}$ and
$p_{j}^{x}=p_{j}^{\bot}\cos\theta,\quad p_{j}^{x}=p_{j}^{\bot}\sin\theta,$ (61)
with $p_{j}^{\perp}=\sqrt{|p_{j}^{x}|^{2}+|p_{j}^{y}|^{2}}$ and $\tan\theta=p_{j}^{x}/p_{j}^{y}$. Then the system can be rewritten in a more compact from by a unitary transformation
$U(z,\theta)=\left(\begin{array}[]{cc}e^{i\theta}&0\\ 0&e^{-i\theta}\end{array}\right),$ (62)
then the system becomes
$\Pi^{\prime}_{j-1}\vec{v}_{j-1}+\Pi^{\prime}{}_{j}^{\dagger}\vec{v}_{j+1}+M_{j}\vec{v}_{j}=E\vec{v}_{j},$ (63)
with the transformed hopping coupling matrix $\Pi^{\prime}_{j}=U(z,\theta/2)\Pi_{j}U^{\dagger}(z,\theta/2)$ being
$\Pi^{\prime}_{j}=p_{j}^{\perp}\sigma_{x}+p_{j}^{z}\sigma_{z},$ (64)
and $\vec{v}_{j}=U(z,\theta)\vec{u}_{j}=\left(\begin{array}[]{cc}v_{j\uparrow}&v_{j\downarrow}\end{array}\right)^{\intercal}$.
Before proving the absence of the MEs when satisfying the criteria, we first identify the chiral symmetry of the system under this condition, which is highly related to the disappearance of the MEs. We notice that under the unitary operator ${\cal O}=\sigma_{y}$, both coupling matrices change sign, i.e.,
${\cal O}\Pi^{\prime}_{j}{\cal O}^{\dagger}=-\Pi^{\prime}_{j},\quad{\cal O}M_{j}{\cal O}^{\dagger}=-M_{j}.$ (65)
Therefore both $\vec{v}_{j}$ and ${\cal O}\vec{v}_{j}$ are the eigenstate of the system, with the corresponding energies come in pair as $(E,-E)$.
With this properties, we proceed to calculate the characteristic polynomial $P(E)$. We first consider the eigen-equation $H\lvert v\rangle=E\lvert v\rangle$ under the twisted boundary condition
$\begin{array}[]{c}e^{ik_{x}}(p_{j-1}^{z}v_{j-1\uparrow}+p_{j-1}^{\perp}v_{j-1\downarrow})+e^{-ik_{x}}(p_{j}^{z*}v_{j+1\uparrow}+p_{j}^{\perp*}v_{j+1\downarrow})+m_{j}v_{j\uparrow}=Ev_{j\uparrow},\\ e^{ik_{x}}(p_{j-1}^{\perp}v_{j-1\uparrow}-p_{j-1}^{z}v_{j-1\downarrow})+e^{-ik_{x}}(p_{j}^{\perp*}v_{j+1\uparrow}-p_{j}^{z*}v_{j+1\downarrow})-m_{j}v_{j\downarrow}=Ev_{j\downarrow},\end{array}$ (66)
then we locally rotate the bases as $\psi_{j}^{\pm}=v_{j\uparrow}\pm iv_{j\downarrow}$, and we can rewrite the eigen-equation as
$\begin{array}[]{c}e^{ik_{x}}(p_{j-1}^{z}+ip_{j-1}^{\perp})\psi_{j-1}^{-}+e^{-ik_{x}}(p_{j}^{z*}+ip_{j}^{\perp*})\psi_{j+1}^{-}+m_{j}\psi_{j}^{-}=E\psi_{j}^{+}\\ e^{ik_{x}}(p_{j-1}^{z}-ip_{j-1}^{\perp})\psi_{j-1}^{+}+e^{-ik_{x}}(p_{j}^{z*}-ip_{j}^{\perp*})\psi_{j+1}^{+}+m_{j}\psi_{j}^{+}=E\psi_{j}^{-}\end{array}$ (67)
By the commensurate approximation, we approximate the irrational number $\alpha$ by the commensurate size, then $L$ becomes the size of the unit cell and the system exhibits the periodic structure. Then to calculate the characteristic polynomial $P(E)=\det|H-EI|$, we can perform the Fourier transform, i.e. we multiply by $e^{i2\pi\alpha jm}$ and sum over $j$, then finding characteristic polynomial $P(E)=\det|H-EI|$ becomes finding the determinant $P(E)=\det|A(E)|$, with $A(E\text{)}$ being
$A=\left(\begin{array}[]{cccc}A_{11}&A_{12}&\dots&A_{L1}\\ A_{21}&\ddots&&\vdots\\ \vdots&&\ddots&\vdots\\ A_{L1}&\dots&\dots&A_{LL}\end{array}\right)_{2L\times 2L},$
by denoting the hopping coefficients $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$, each block being a 2-by-2 matrix as
$A_{mj}=e^{i2\pi\alpha jm}\left(\begin{array}[]{cc}t_{j-1}^{+}e^{-i(2\pi\alpha m+k_{x})}+t_{j}^{-*}e^{i(2\pi\alpha m+k_{x})}+m_{j}&-E\\ -E&t_{j-1}^{-}e^{-i(2\pi\alpha m+k_{x})}+t_{j}^{+*}e^{i(2\pi\alpha m+k_{x})}+m_{j}\end{array}\right)$ (68)
This form has the advantages that within each element, the energy term $E$ is decoupled from the hopping $p_{j}$ and on-site term $m_{j}$, facilitating the discussion.
To prove that in this case, the system exhibits pure phase with MEs, we are going to prove that the effective coefficients of the dispersion as the we iterating the commensurate approximated system size $L$ will saturate to a energy-independent parameter. First of all, due to the presence of the chiral symmetry, the spectrum of the system is symmetric in energy $E$ with respective to zero energy. Therefore the characteristic polynomial $P(E)=\det|A|$ can only be the polynomial of energy $E$ with even power. In addition, due to the periodic structure of the Hamiltonian, the characteristic polynomial is the function of $Lk_{x}$ and $Lk_{y}$. Now we consider the case of odd $L$.
We first prove the coefficients that give rise to extended and localized states are energy independent. We first investigate the parameters attached to the leading dispersion $\cos Lk_{x}$ and $\cos Lk_{y}$. Since each term within the determinant is obtained by a multiply of $2L$ terms, the terms consisting $\cos Lk_{x}$ comes from the multiplication of $L$ folds $e^{ik_{x}}$ and $L$ folds energies $E$. This indicates that such terms are illegal under the chiral symmetry since such term is of odd power of energy $\sim E^{L}\cos Lk_{x}$, which is also the same for $\cos k_{y}$. So the parameters associated with dispersion $\cos Lk_{x}$ and $\cos Lk_{y}$ vanish. Then the coefficients that give rise to extended and localized states, become the parameters along with the dispersion $\cos 2Lk_{x}$ and $\cos 2Lk_{y}$. These two terms can only be obtained by multiply all the diagonal terms in $\det A$, therefore they are energy independent.
Now we prove the parameters associated with the critical states, which are linked to the dispersion $\cos n_{x}Lk_{x}\cos n_{y}Lk_{y}$ with $n_{x},n_{y}\geq 1\in\mathbf{\mathbb{Z}}$, are also energy independent. Since we are considering the case $p_{j}^{x}/p_{j}^{y}=\mathrm{Const.}\in\mathbb{R}$ and $M_{j}$ are purely quasiperiodic, there ate basically two choices for the form of the hopping coupling matrix $\Pi_{j}$: One is that the hopping coupling matrix $\Pi_{j}$ are uniform while the on-site matrix $M_{j}$ are purely quasiperiodic. And the other choice is that both $\Pi_{j}$ and $M_{j}$ are purely quasiperiodic.
We first prove that for the case $\Pi_{j}$ are uniform, the coefficient associated with $\cos Lk_{x}\cos Lk_{y}$ is energy independent. This is because the dispersion $\cos Lk_{x}\cos Lk_{y}$ comes from the product of combination of $L$ folds $p_{j}$ terms and combination of $L$ folds $m_{j}$ terms, which contribute to $L$ folds $e^{ik_{x}}$ and $e^{ik_{y}}$, respectively. Then the total multiplication of $2L$ elements contribute the $\cos Lk_{x}\cos Lk_{y}$, leaving no chance for involving $E$.
Then we consider the case both $\Pi_{j}$ and $M_{j}$ are purely quasiperiodic, in which case the leading dispersion is $\cos 2Lk_{x}\cos Lk_{y}$, $\cos Lk_{x}\cos 2Lk_{y}$ or $\cos 2Lk_{x}\cos 2Lk_{y}$, whose coefficients are all energy independent. A key ingredient is that the hopping coupling matrix exhibits the off-diagonal modulation $(V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.})$, which makes the $p_{j}$ contributes the $k_{x}$ and $k_{y}$ simultaneously. Therefore, if $\Pi_{j}$ are quasiperiodic, $\cos Lk_{x}\cos Lk_{y}$ dispersion will disappear. The reason is that $\cos Lk_{x}\cos Lk_{y}$ comes from the multiplication of $L$ folds $p_{j}$ terms and $L$ folds $E$ terms, which is not allowed under the chiral symmetry. So the leading order contributing to the critical orbital in this case is $\cos Lk_{x}\cos 2Lk_{y}$ or $\cos 2Lk_{x}\cos Lk_{y}$. The first case corresponds multiplication of $L$ folds $p_{j}$ terms and $L$ folds $m_{j}$ terms, which are in total $2L$ terms and leaves no chance to involve the extra terms. So the coefficients associated with $\cos Lk_{x}\cos Lk_{y}$ are energy independent. Similarly, the $\cos 2Lk_{x}\cos Lk_{y}$ originates from the product of $2L$ folds $p_{j}$ terms. Generically this will lead to $\cos 2Lk_{x}\cos Lk_{y}$ and $\cos 2Lk_{x}\cos 2Lk_{y}$, and both of them are energy independent. It worthwhile to note that he former one comes from the destruction of relative phases, which gives rise to the $\cos 2Lk_{x}\cos Lk_{y}$.
Appendix B Exact solvability of QP systems from local constraint
Now we show the exact solvability for the spinful quasiperiodic system when the hopping coupling matrices satisfy the condition
$\det|\Pi_{j}|=0,\quad p_{j}^{s}=\mathrm{Const.},$ (69)
with off-diagonal terms of $M_{j}$ being purely constant or purely quasiperiodic.
We first consider the constraint for spin-conserved process, in which case the hopping coupling matrix after the local unitary transformation is given by
$\tilde{\Pi}_{j}=\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right).$ (70)
Without losing generality, we set $t=1$ here, then the eigenvalue equations $H\lvert\Psi\rangle=E\lvert\Psi\rangle$ with generic on-site coupling matrix $M_{j}$ become
$\begin{split}u_{j-1\uparrow}+u_{j+1\uparrow}+M_{j}^{11}u_{j\uparrow}+M_{j}^{12}u_{j\downarrow}=Eu_{j\uparrow},\\ M_{j}^{21}u_{j\uparrow}+M_{j}^{22}u_{j\downarrow}=Eu_{j\downarrow}.\end{split}$ (71)
Notice that here $u_{j\uparrow}$ and $u_{j\downarrow}$ refer to the coefficients of the wavefunction of the dressed particle after the unitary transformation. We can then obtain the reduced eigenvalue equation
$u_{j-1\uparrow}+u_{j+1\uparrow}+\Big{(}M_{j}^{11}+\frac{|M_{j}^{12}|^{2}}{E-M_{j}^{22}}\Big{)}u_{j\uparrow}=Eu_{j\uparrow},$ (72)
which describe the effective motion with nearest neighbor hopping and the effective on-site potential is energy dependently dressed by the local modulation from the accompanied internal degree of freedom.
From this expression, we can also see the reason why we require the off-diagonal terms of $M_{j}$ being purely QP or purely constant. If the off-diagonal terms of $M_{j}$ contain a constant plus a QP term, then its square will contain both a single QP frequency and double QP frequency, which generically is not analytically solvable.
We then investigate the constraint for spin-conserved process, in which case the hopping coupling matrix after the local unitary transformation is given by
$\tilde{\Pi}_{j}=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right),$ (73)
Again we we set $\mu=1$ here for simplicity, then the eigenvalue equations $H\lvert\psi\rangle=E\lvert\psi\rangle$ with generic on-site coupling matrix $M_{j}$ become
$\begin{split}u_{j-1\downarrow}+M_{j}^{11}u_{j\uparrow}+M_{j}^{12}u_{j\downarrow}=Eu_{j\uparrow},\\ u_{j+1\uparrow}+M_{j}^{21}u_{j\uparrow}+M_{j}^{22}u_{j\downarrow}=Eu_{j\downarrow}.\end{split}$ (74)
Then the effective eigen-equation for one for the species is
$\frac{M_{j-1}^{21}}{E-M_{j-1}^{22}}u_{j-1\uparrow}+\frac{M_{j}^{12}}{E-M_{j}^{22}}u_{j+1\uparrow}+\Big{(}M_{j}^{11}+\frac{1}{E-M_{j-1}^{22}}+\frac{|M_{j}^{12}|^{2}}{E-M_{j}^{22}}\Big{)}u_{j\uparrow}=Eu_{j\uparrow},$ (75)
which describes the the effective hopping and on-site potentials are energy dependently dressed by the transfer and local modulation of the accompanied internal degree of freedom.
For the combination of both are similar, where the hopping coupling matrix is given by
$\tilde{\Pi}_{j}=\left(\begin{array}[]{cc}1&1\\ 0&0\end{array}\right),$ (76)
Similarly we we set both $t=1$ and $\mu=1$, then the eigenvalue equations $H\lvert\psi\rangle=E\lvert\psi\rangle$ with generic on-site coupling matrix $M_{j}$ become
$\begin{split}u_{j+1\uparrow}+u_{j+1\downarrow}+u_{j-1\uparrow}+M_{j}^{11}u_{j\uparrow}+M_{j}^{12}u_{j\downarrow}=Eu_{j\uparrow},\\ u_{j-1\uparrow}+M_{j}^{21}u_{j\uparrow}+M_{j}^{22}u_{j\downarrow}=Eu_{j\downarrow}.\end{split}$ (77)
And the effective eigen-equation becomes
$\Big{(}1+\frac{M_{j}^{12}}{E-M_{j}^{22}}\Big{)}u_{j-1\uparrow}+\Big{(}1+\frac{M_{j+1}^{21}}{E-M_{j+1}^{22}}\Big{)}u_{j+1\uparrow}+\Big{(}M_{j}^{11}+\frac{1}{E-M_{j+1}^{22}}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}\Big{)}u_{j\uparrow}=Eu_{j\uparrow}.$ (78)
Appendix C 1D QP mosaic lattice model
In this section, we introduce the one-dimensional quasiperiodic mosaic lattice model Wang et al. (2020a); Zhou et al. (2023) and rewrite it into our 1D spinful QP framework. The generic mosaic lattice refers that the modulation of on-site potential and hopping coefficients are in a mosaic manner, namely the the QP modulation or the uniform modulation only appears every $\kappa$ sites, as we will see in the following.
The original type-I and type-II mosaic lattice model is distinguished by the presence of alternating mosaic pattern in hopping coefficients. If the QP modulation only exhibits in the on-site potential and the hopping coefficients are always uniform or always quasiperiodic, then the mosaic models belongs to type-I mosaic model. The Hamiltonian of the type-I mosaic lattice model is given by Wang et al. (2020a):
$H=\lambda\sum_{j}(c_{j}^{\dagger}c_{j+1}+{\rm h.c.})+2\sum_{j}V_{j}c_{j}^{\dagger}c_{j},$ (79)
with the on-site potential given by
$V_{j}=\begin{cases}V\cos(2\pi\alpha j+\theta),&j=0\quad{\rm mod}\leavevmode\nobreak\ \kappa,\\ 0,&j\neq 0\quad{\rm mod}\leavevmode\nobreak\ \kappa,\end{cases}$ (80)
where $c_{j}$ is the annihilation operator at site $j$, and $\lambda$, $V$, and $\theta$ denote the nearest-neighbor hopping coefficient, on-site potential amplitude, and phase offset, respectively. $\alpha$ is an irrational number, and $\kappa$ is an integer. The quasiperiodic potential periodically occurs every $\kappa$ sites.
If the system exhibits mosaic modulation in hopping coefficients, then it belongs to type-II mosaic model Zhou et al. (2023) and the Hamiltonian is given by
$H=\sum_{j}(t_{j}c_{j}^{\dagger}c_{j+1}+\text{H.c.})+\sum_{j}V_{j}c_{j}^{\dagger}c_{j},$ (81)
here both the quasiperiodic hopping coefficient $t_{j}$ and on-site potential $V_{j}$ are in mosaic manners, with the hopping coefficient given by
$t_{j}=\begin{cases}2t\cos(2\pi\alpha j+\theta),&j=0\bmod\kappa,\\ \lambda,&j=1\bmod\kappa,\\ \lambda,&\text{otherwise},\end{cases}$ (82)
and the on-site potential given by
$V_{j}=\begin{cases}2t\cos(2\pi\alpha j+\theta),&j=0\bmod\kappa,\\ 2t\cos[2\pi\alpha(j-1)+\theta],&j=1\bmod\kappa,\\ 0,&\text{otherwise}.\end{cases}$ (83)
If we focus on the case $\kappa=2$, then both two moasic lattice models exhibit a even-odd sublattice structure, which can be recasted into the spinful QP chain model. We can relabel the even and odd site as spin-up and spin-down degrees of freedom and further relabel the site index, then the Hamiltonian can be rewritten as
$H_{\mathrm{M}}=\sum_{j}\lambda(c^{\dagger}_{j+1}\sigma_{-}c_{j}+\mathrm{h.c.})+\sum_{j}c_{j}^{\dagger}V_{j}^{\mathrm{M}}c_{j},$ (84)
where $c_{j}=(c_{j,\uparrow},c_{j,\downarrow})^{\intercal}$ is the spinor for the annihilation operators. The parameter $V_{j}^{\mathrm{M}}$ controls the type-I and type-II mosaic lattice, which respectively given by
$V_{j}^{\mathrm{M,I}}$ $=$ $2V_{0}V_{j}^{\mathrm{d}}\Lambda_{+}+\lambda\sigma_{x},$ (85)
$V_{j}^{\mathrm{M,II}}$ $=$ $2tV_{j}^{\mathrm{d}}(\sigma_{0}+\sigma_{x}).$ (86)
Appendix D 1D spinless QP model
In this section, we show that the 1D spinless QP model with nearest neighbour hopping can also be unified within the spinful QP systems, under the Majorana representation. For a generic 1D spinless QP chain model, the tight binding Hamiltonian is given by
$H=t_{j}\sum_{j}(c_{j}^{\dagger}c_{j+1}+\mathrm{h.c.})+\sum_{j}V_{j}c_{j}^{\dagger}c_{j}.$ (87)
The $t_{j}$ and $V_{j}$ are the hopping coefficients and on-site potential, respectively. Such 1D spinless QP chain can be recasted into the spinful QP chain by introducing the Majorana representation
$c_{j}=\frac{1}{2}(\gamma_{B,j}+i\gamma_{A,j}),\quad\quad c_{j}^{\dagger}=\frac{1}{2}(\gamma_{B,j}-i\gamma_{A,j}),$ (88)
with $\gamma_{\sigma,j}=\gamma_{\sigma,j}^{\dagger}$ and $\{\gamma_{\sigma,j},\gamma_{\sigma;,j^{\prime}}\}=2\delta_{\sigma\sigma^{\prime}}\delta_{jj^{\prime}}$. In this basis, the Hamiltonian becomes
$H$ $=$ $\frac{t_{j}}{2}\sum_{j}(i\gamma_{B,j}\gamma_{A,j+1}+i\gamma_{B,j+1}\gamma_{A,j}+)$ (89)
$+$ $\sum_{j}\frac{V_{j}}{2}(1+i\gamma_{B,j}\gamma_{A,j}).$
One can neglect the last summation, which leads to a constant, then the generic Hamiltonian is unified within our 1D spinful formalism, with the hopping coupling matrix and the on-site matrix given by
$\Pi_{j}=\frac{t_{j}}{4}\sigma_{y},\quad M_{j}=\frac{V_{j}}{4}\sigma_{y}.$ (90)
One can notice that, its characteristic is that both hopping coupling and on-site matrix shares the same Pauli component $\sigma_{y}$. This formalism unifies all the well known 1D spinless QP model in this context, including the AA model Aubry and André (1980), the extended AA (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015), the generalized AA (GAA) model Ganeshan et al. (2015), or the GPD model in some context. These are summarized in the table. 3.
Parameters Models
$t_{j}=t\quad V_{j}=2VV_{j}^{\mathrm{d}}$ AA model
$t_{j}=(t+\mu)V_{j}^{\mathrm{od}}\quad V_{j}=2V_{0}V_{j}^{\mathrm{d}}$ EAA model
$t_{j}=t\quad V_{j}=2V_{0}V_{j}^{\mathrm{d}}/(1-aV_{j}^{\mathrm{d}})$ GAA model
Table 3: The parameters of different 1D spinless QP model
Figure 7: The proposed laser configuration enables the realization of a model featuring all three fundamental types of mobility edges. Laser beams ${\bf E}_{1}$ and ${\bf E}_{2}$ with wavelengths $\lambda_{1}$ and $\lambda_{2}$ ,individually generate spin-independent lattices, while ${\bf E}_{3}$ and ${\bf E}_{4}$ operating at ‘tune-out’ wavelength $\lambda_{3}$ form a periodic Raman potential $(\,^{87}\text{Rb atoms})$.
In this setup, the ratio $\alpha=k_{s}/k_{p}\approx 0.83$, which is experimentally feasible for observing MEs.
Appendix E Experimental Scheme
In this section, we provide a more detailed description of the experimental scheme for realizing the Hamiltonian Eq. (51) in the main text. Without loss of generality, we consider ||$\ket{\uparrow}\equiv\ket{F=1,m_{F}=-1}$ and ||$\ket{\downarrow}\equiv\ket{1,0}$ of ${}^{87}{\rm Rb}$ atoms as an example, while all results remain applicable to other alkali atoms. After applying the unitary transformation $\sigma_{z}\rightarrow\sigma_{x}$ and $\sigma_{x}\rightarrow-\sigma_{z}$, the target Hamiltonian can be rewritten as
$H=\left[\frac{p_{z}^{2}}{2m}+{\cal V}_{s}(z)\right]\otimes\sigma_{0}+{\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2-M_{0}\sigma_{z},$ (91)
where ${\cal V}_{p}(z)$ and ${\cal V}_{s}(z)$ represent the primary and secondary optical lattices, respectively. These lattices are generated by two sets of standing waves and a periodic Raman potential. To construct the secondary lattice, we require a standing wave of the form
${\bf E}_{1}=2E_{1}\hat{e}_{x}e^{{\rm i}(\phi_{1}+\phi_{1}^{\prime}/2)}\cos(k_{1}z-\phi_{1}^{\prime}/2),$ (92)
where $\phi_{1}$ (without the prime) is the initial phase of the incident beam, and $\phi_{1}^{\prime}$ is an additional phase acquired by ${\bf E}_{1}$ before being reflected back to the atoms. Given that the modulus of the wave vector is $k_{1}=2\pi/\lambda_{1}$, and ignoring the overall constant energy shift, the standing wave field ${\bf E}_{1}$ induces a spin-independent potential given by
${\cal V}_{1}(z)$ $=\frac{V_{1}}{2}\cos(2k_{1}z-\phi_{1}^{\prime}),$
$\text{where}\quad V_{1}$ $=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,10}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (93)
In the above, the transition matrix elements are $\Omega_{F^{\prime},m_{F}^{\prime};\sigma,nq}^{(J)}=-\langle F^{\prime},m_{F}^{\prime}|{\bf d}|\sigma\rangle E_{n}/\hbar$ and $\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}$ are detunings of laser field from different levels. Here, $J=1/2,3/2$ correspond to the $D_{1}$ and $D_{2}$ lines respectively, and $\sigma=\uparrow,\downarrow$ labels the spin states. The index $n$ denotes the laser beam, while $q=0,\pm$ represents the polarization of laser field. For a laser wavelength of $\lambda_{1}\approx 1064\,{\rm nm}$, the overall optical shifts induced by the linearly polarized ($q=0$) field from ${\bf E}_{1}$ are nearly identical for the two spin states. Comparing this with Eq. (51), we find
$k_{s}=2k_{1},\quad V_{s}=V_{1}/2,\quad\phi_{s}=-\phi_{1}^{\prime}.$ (94)
On the other hand, the primary lattice ${\cal V}_{p}(z)$ is introduced through three laser fields, given by
${\bf E}_{2}$ $=2E_{2}\hat{e}_{x}e^{{\rm i}(\phi_{2}+\frac{\phi_{2}^{\prime}}{2})}\cos(k_{2}z-\frac{\phi_{2}^{\prime}}{2}),$
${\bf E}_{3}$ $=2E_{3}\hat{e}_{x}e^{{\rm i}(\phi_{3}+\frac{\phi_{3}^{\prime}}{2})}\cos(k_{3}z-\frac{\phi_{3}^{\prime}}{2}),$
${\bf E}_{4}$ $=E_{4}\hat{e}_{z}e^{-{\rm i}k_{3}y+{\rm i}\phi_{4}}.$ (95)
We first define the spin-independent part of the primary lattice ${\cal V}_{p}(z)$ as a periodic potential generated by the standing wave ${\bf E}_{2}$, given by
${\cal V}_{p}(z)$ $=\frac{V_{2}}{2}\cos(2k_{2}z-\phi_{2}^{\prime}),$
$\text{where}\quad V_{2}$ $=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,20}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (96)
Then the standing wave ${\bf E}_{3}$ and the traveling wave ${\bf E}_{4}$ together generate a periodic Raman potential along the z-direction:
${\cal V}_{p}(z)$ $=M_{34}e^{{\rm i}(\phi_{4}-\phi_{3}-\phi_{3}^{\prime}/2)}\cos(k_{3}z-\phi_{3}^{\prime}/2),$
$\text{where}\quad M_{34}$ $=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\Omega_{F^{\prime},m_{F}^{\prime};\uparrow,30}^{(J)*}\Omega_{F^{\prime},m_{F}^{\prime};\downarrow,4-}^{(J)}}{\sqrt{2}\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (97)
Here, we adopt the following gauge convention for the spherical basis: $\hat{e}_{z}=-1(\hat{e}_{+}-\hat{e}_{-})/\sqrt{2}$ and $\hat{e}_{y}={\rm i}(\hat{e}_{+}+\hat{e}_{-})/\sqrt{2}$. For simplicity, we choose ‘tune-out’ wavelength $\lambda_{3}\approx 790\,{\rm nm}$ Wen et al. (2021). As a result, the lattice depth generated by ${\bf E}_{3}$ is nealy zero, since the optical shifts from the $D_{1}$ and $D_{2}$ lines cancel each other out footnote_Exp . By comparing this expression with Eq. (E), the following identities hold:
$k_{3}=2k_{2},\quad M_{34}=V_{2}/2,\quad\phi_{3}^{\prime}/2=\phi_{2}^{\prime}.$ (98)
The exponential phase factor $\phi_{4}-\phi_{3}-\phi_{2}^{\prime}=0$ can be stabilized by locking the relative phase $\phi_{4}-\phi_{3}$ using a phase-locking technique Jotzu et al. (2014); Görg et al. (2019); Wang et al. (2021c). The two-photon detuning of Raman coupling is defined as $\delta/2=(\Delta E+\omega_{3}-\omega_{4})/2\equiv M_{0}$, where $\Delta E$ is the energy splitting between $\ket{\uparrow}$ and $\ket{\downarrow}$, typically on the order of tens of MHz and $\omega_{3,4}$ denote the corresponding frequencies of ${\bf E}_{3,4}$. The sources for ${\bf E}_{2}$ and ${\bf E}_{3,4}$ can be generated using a Raman fiber laser with second-harmonic generation (SHG) cavity, which provides suitable wavelengths of both $\lambda_{2}$ and $\lambda_{3,4}$ at the same time. The additional phase $\phi_{3}^{\prime}$ and $\phi_{2}^{\prime}$ acquired along the reflective path can be finely tuned by varying the optical path length of the reflected beam.
Finally, we see that the QD lattice in Eq. (51) can be realized using the proposed scheme when the following conditions are satisfied:
$k_{p}=2k_{2},\quad V_{p}=V_{2}/2,\quad\phi_{p}=-\phi_{2}^{\prime}.$ (99)
This experimental scheme leads to a ratio of $\alpha=k_{p}/k_{s}\approx 0.67$. The nearest-neighbor hopping amplitude $t$, determined by the lattice potential $V_{0}$, approximately follows the relation $t\approx 4V_{0}^{3/4}e^{-2\sqrt{V_{0}}}/\sqrt{\pi}$ Li et al. (2017). To suppress the nearest-neighbor hopping for spin-down atoms, the parameters $V_{s}=2.5,\,V_{p}=10,\,M_{0}=1.5$ are chosen such that the hopping amplitude $t$ for spin-up atoms is about one order of magnitude larger than that for spin-down atoms.